### v1

# Uniqueness of asymptotically cylindrical steady gradient Ricci solitons

Abstract.
We show that the Bryant soliton is the unique asymptotically cylindrical steady gradient Ricci soliton, in any dimension $n\geq 3$ and without any curvature assumptions. This generalizes a celebrated theorem of Brendle. We also prove that any steady gradient Ricci soliton asymptotic to a cylinder over the homogeneous lens space $\mathbb{S}^{2m+1}/\mathbb{Z}_{k}=L_{m,k}$, for $m\geq 1$ and $k\geq 3$, is a noncollapsed Appleton soliton on the complex line bundle $\mathcal{O}(-k)$ over $\mathbb{CP}^{m}$. In dimension 4, our results lead to a classification of steady gradient Ricci soliton singularity models on smooth manifolds which possess a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$.
1. Introduction
A Ricci soliton is a complete $n$-dimensional Riemannian manifold $(M^{n},g)$ with a complete vector field $X$ such that
(1.1) $\operatorname{Ric}=\frac{1}{2}\mathcal{L}_{X}g+\lambda g,$
where $\mathcal{L}$ is the Lie derivative, and $\lambda\in\mathbb{R}$ determines the soliton type: shrinking ($\lambda>0$), steady ($\lambda=0$), or expanding ($\lambda<0$). If $X=\nabla f$ for a function $f:M\to\mathbb{R}$, then $(M^{n},g,f)$ is called a gradient Ricci soliton and the equation (1.1) becomes
(1.2) $\operatorname{Ric}=\mathrm{Hess}_{f}+\lambda g.$
Throughout this article, we abbreviate GRS = gradient Ricci soliton.
Ricci solitons are self-similar solutions of the Ricci flow. Shrinking and steady Ricci solitons are special instances of ancient solutions, which are solutions of Ricci flow defined for all negative time. Ancient solutions arise as blowup limits of Ricci flow at singularities; we call these blowup limits singularity models. Classifying Ricci solitons and ancient solutions provides a way to understand singularity models, which is in turn crucial for understanding singularity formation in Ricci flow.
In dimension 3, Perelman [perelman1] showed that finite-time singularity models of Ricci flow are ancient $\kappa$-solutions: non-flat, $\kappa$-noncollapsed ancient solutions with bounded and nonnegative sectional curvature. The classification of ancient $\kappa$-solutions in dimension 3 was completed by Brendle [brendle-acta], building on techniques developed in his work on 3-dimensional steady GRSs [brendle2013rotational] and the foundations laid by many other authors. For more information on Ricci solitons and ancient solutions in dimensions $\leq 3$, we refer to the books [chow2023ricci, chowluni] and references therein.
In dimensions $\geq 4$, ancient $\kappa$-solutions no longer capture all singularity models of Ricci flow. In fact, singularity models in general need not enjoy any curvature bound unless one imposes strong positivity conditions on the initial data, such as positive isotropic curvature in dimension 4 [hamilton-pic]. Yet, all known classification results for Ricci solitons and ancient solutions in dimensions $\geq 4$ impose curvature hypotheses alongside other structural assumptions; see for instance [brendle2014rotational, brendle-naff, zz, deng-zhu, chodosh, chodosh-fong, cds]. This raises the natural question: can one classify Ricci solitons and/or ancient solutions in dimensions $\geq 4$ without imposing any curvature assumptions?
In this paper, we answer this question affirmatively for steady GRSs that are asymptotic to cylinders over certain spherical space forms. The notion of asymptotic cylindricity is standard and arises naturally in various situations; see e.g. [brendle2013rotational, deng-zhu, bamler-3].
Definition 1.1.
Let $n\geq 3$. Let $\mathcal{S}=\mathbb{S}^{n-1}/\Gamma$ be a spherical space form of dimension $n-1$. A steady GRS $(M^{n},g,f)$ is said to be asymptotically $\mathcal{S}$-cylindrical if the following holds. Let $\Phi_{t}$ be the flow of the vector field $\nabla f$. Take an arbitrary sequence $p_{m}$ of marked points going to infinity and consider the rescaled flows
(1.3) $\hat{g}^{(m)}(t)=r_{m}^{-1}\Phi_{r_{m}t}^{*}g,$
with $r_{m}R(p_{m})=\frac{n-1}{2}+o(1)$, where $R(p_{m})$ is the scalar curvature at $p_{m}$. We require that the flows $(M,\hat{g}^{(m)}(t),p_{m})$ converge as $m\to\infty$ (without passing to a subsequence) in the Cheeger–Gromov sense to a family of shrinking cylinders $(\mathcal{S}\times\mathbb{R},\bar{g}(t))$, $t\in(0,1)$. The metric $\bar{g}(t)$ is given by
(1.4) $\bar{g}(t)=(n-2)(2-2t)\bar{g}_{\mathcal{S}}+dz\otimes dz,$
where $\bar{g}_{\mathcal{S}}$ is a metric on $\mathcal{S}$ with constant sectional curvature 1.
Definition 1.2.
We call $(M^{n},g,f)$ asymptotically cylindrical if it is asymptotically $\mathbb{S}^{n-1}$-cylindrical.
Remark 1.3.
In [brendle2014rotational], Brendle also required the scalar curvature to satisfy $\frac{\Lambda_{1}}{r}\leq R\leq\frac{\Lambda_{2}}{r}$, where $\Lambda_{1},\Lambda_{2}>0$ and $r$ is the distance to a fixed point in $M$. It has since been shown that Definition 1.1 implies this; see [deng-zhu]*Lemma 6.5 and [CDN]*§3.
We now describe our results.
1.1. Uniqueness of the Bryant soliton
A famous theorem of Brendle [brendle2013rotational, brendle2014rotational] states that any asymptotically cylindrical steady GRS of dimension $n\geq 3$ with positive sectional curvature must be rotationally symmetric, hence isometric to the Bryant soliton [bryant]. In three dimensions, the hypotheses are met for any non-flat, $\kappa$-noncollapsed steady GRS, and Brendle’s result in this setting resolved a major conjecture left open in Perelman’s first paper [perelman1]. Ten years later, Zhao and Zhu [zz] generalized Brendle’s theorem in dimensions $\geq 4$ by relaxing the positive sectional curvature condition to the weaker condition $2P<R$, where $P$ is the operator norm of the Riemann curvature acting on symmetric 2-tensors.
Our first result proves uniqueness of the Bryant soliton in a much broader setting. We show that uniqueness holds in any dimension only assuming asymptotic cylindricity, without any curvature hypotheses.
Theorem 1.4.
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$. Then $(M,g,f)$ is the Bryant soliton (i.e. they are isometric up to scale).
In dimension 4, Deng and Zhu [deng-zhu]*Theorem 1.5 established a sufficient condition for a steady GRS to be asymptotically cylindrical. Combining Theorem 1.4 with their result, we get:
Corollary 1.5.
Any 4-dimensional $\kappa$-noncollapsed steady GRS with positive Ricci curvature and linear curvature decay must be the Bryant soliton.
This affirmatively answers a conjecture posed by Deng and Zhu in the same paper. Corollary 1.5 is comparable to Brendle’s 4-dimensional result [brendle2014rotational], which reaches the same conclusion assuming positive isotropic curvature instead of ‘positive Ricci curvature and linear curvature decay’.
Theorem 1.4 is proved using Brendle’s Killing field technique, which upgrades asymptotic symmetries to exact ones then exploits them for rigidity and uniqueness. We generalize this technique to hold on any asymptotically $\mathcal{S}$-cylindrical steady GRS, without the curvature contraints imposed in [brendle2013rotational, brendle2014rotational, zz]. Although every such GRS has positive curvature operator outside a compact set [deng-zhu] (see also [bamler-3]*Proposition 3.24), one cannot directly generalize the arguments in [brendle2013rotational, brendle2014rotational, zz] as they crucially rely on global curvature conditions. See §1.6 for an outline of our proofs, and [chodosh, chodosh-fong, brendle-acta, brendle-naff] for other applications of Brendle’s Killing field technique.
The generality of our methods leads to the non-existence of asymptotically $\mathcal{S}$-cylindrical steady GRSs for certain $\mathcal{S}$. When $\mathcal{S}=\mathbb{RP}^{n-1}$, non-existence follows from minor modifications to the proof of Theorem 1.4.
Theorem 1.6.
There is no asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS for any $n\geq 3$.
1.2. Uniqueness of Appleton’s noncollapsed solitons
In [appleton], Appleton discovered a new family of $\kappa$-noncollapsed steady GRSs in each even dimension $\geq 4$, and indicated numerical evidence of these solitons arising as Ricci flow singularity models. A uniqueness theorem is therefore in order, and this is our next result. To state this, we first review the construction of Appleton’s solitons.
For $m,k\geq 1$, the principal $U(1)$ Hopf fibration $\pi:\mathbb{S}^{2m+1}\to\mathbb{CP}^{m}$ induces a free action of the cyclic subgroup $\mathbb{Z}_{k}\subset U(1)$ on $\mathbb{S}^{2m+1}$. Equip $\mathbb{S}^{2m+1}$ with a Berger metric $\tilde{g}_{a,b}$ for some $a,b>0$, that is
(1.5) $\tilde{g}_{a,b}=a^{2}\sigma\otimes\sigma+b^{2}\pi^{*}\bar{g}_{\mathbb{CP}^{m}}$
where $\sigma$ is a 1-form dual to the Hopf fibers, and $\bar{g}_{\mathbb{CP}^{m}}$ is the Fubini–Study metric with holomorphic sectional curvature 4. Then the action of $\mathbb{Z}_{k}$ is isometric on $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$, and the quotient is the lens space $(L_{m,k},g_{a,b})$. When $a=b=\kappa$, the metrics $\tilde{g}_{a,b}$ and $g_{a,b}$ have constant sectional curvature $\kappa^{-2}$.
Appleton considered metrics on $(0,\infty)\times L_{m,k}$ of the form
(1.6) $g=ds^{2}+g_{a(s),b(s)}$
which are smoothly completed at $s=0$ by inserting a central $\mathbb{CP}^{m}$. The resulting manifold $M$ has dimension $22m+$ 2 and is diffeomorphic to the total space of the complex line bundle $\mathcal{O}_{\mathbb{CP}^{m}}(-k)$. Appleton showed that $(M,g,f)$ defines a steady GRS with $f=f(s)$ whenever $(f,a,b):[0,\infty)\to\mathbb{R}^{3}$ solves the system [^1] Our convention for a steady GRS is $\operatorname{Ric}=\mathrm{Hess}_{f}$, so our $f$ differs from that in [appleton] by a sign.
(1.7) $f^{\prime\prime}$ $=-\frac{a^{\prime\prime}}{a}-2m\frac{b^{\prime\prime}}{b},$
(1.8) $a^{\prime\prime}$ $=2m\left(\frac{a^{3}}{b^{4}}-\frac{a^{\prime}b^{\prime}}{b}\right)-a^{\prime}f^{\prime},$
(1.9) $b^{\prime\prime}$ $=\frac{2m+2}{b}-2\frac{a^{2}}{b^{3}}-\frac{a^{\prime}b^{\prime}}{a}-(2m-1)\frac{(b^{\prime})^{2}}{b}-b^{\prime}f^{\prime},$
with boundary conditions
(1.10) $\begin{gathered}f(0)=0,\quad f^{\prime}(0)=0,\quad f^{\prime\prime}(0)>0,\\ a(0)=0,\quad a^{\prime}(0)=k,\quad b(0)=1,\quad b^{\prime}(0)=0.\end{gathered}$
For each $m,k\geq 1$, define the class of solutions
[…]
We will often write $(M,g,f)\in\mathcal{A}_{m,k}$, even though strictly speaking $(M,g,f)$ is the steady GRS defined by a triple $(f,a,b)\in\mathcal{A}_{m,k}$. Appleton proved the following:
Theorem 1.7 ([appleton]*Theorem 1.3).
Let $m,k\geq 1$.
  1. (a)
    Each $(M,g,f)\in\mathcal{A}_{m,k}$ is $U(m+1)$-invariant, $\kappa$-noncollapsed, and asymptotically $L_{m,k}$-cylindrical (see Appendix A).
  2. (b)
    If $m\geq 1$ and $k\geq m+2$, then $\mathcal{A}_{m,k}$ is nonempty.
As a converse to part (a) and as an analog of Theorem 1.4, we prove the following uniqueness theorem for Appleton’s solitons.
Theorem 1.8.
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$. Then $(M,g,f)\in\mathcal{A}_{m,k}$.
Remark 1.9.
Appleton also conjectured that each $\mathcal{A}_{m,k}$ contains a unique soliton up to scaling and isometry. We do not address this uniqueness question here.
1.3. 4D steady GRSs with homogeneous 3-cylindrical tangent flows
Any 3-dimensional spherical space form that is homogeneous with the round metric (i.e. its isometry group acts transitively) is of the form $SU(2)/\Gamma$ for some finite subgroup $\Gamma\subset SU(2)$ acting from the right. Every finite subgroup of $SU(2)$ is one of the following up to conjugacy (see e.g. [cmp-su2]):
  • The cyclic group $\mathbb{Z}_{k}$ for $k\geq 1$;
  • The binary dihedral group $\tilde{D}_{k}$ of order $4k$ for $k\geq 2$;
  • The binary tetrahedral group $\tilde{T}$ of order 24;
  • The binary octahedral group $\tilde{O}$ of order 48;
  • The binary icosahedral group $\tilde{I}$ of order 120.
We have $L_{1,k}\cong SU(2)/\mathbb{Z}_{k}$. Thus, Theorems 1.4, 1.6, and 1.8 together classify the asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical steady GRSs for $k\geq 1$. To complete the picture, we classify the asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for all other $\Gamma$. For $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, no such solitons exist:
Theorem 1.10.
There is no asymptotically $SU(2)/\Gamma$-cylindrical steady GRS for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
For $\Gamma=\tilde{D}_{k}$ ($k\geq 2$), the story is different. Let $E_{k}$ be the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. The sphere bundles of $E_{k}$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$, and there is a double covering $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$. We will see that any Appleton soliton $A\in\mathcal{A}_{1,2k}$ descends through the map $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$, giving an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS on $E_{k}$ denoted by $A/\mathbb{Z}_{2}$. Moreover, we show that $A/\mathbb{Z}_{2}$ are the only steady GRSs with such asymptotics.
Theorem 1.11.
Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$. Then $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
Combining the 4-dimensional versions of our results (Theorems 1.4, 1.6, 1.8, 1.10, and 1.11), we arrive at the following classification.
Corollary 1.12.
Let $(M^{4},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS, where $\mathcal{S}=SU(2)/\Gamma$ is a homogeneous 3-dimensional spherical space form with $\Gamma\subset SU(2)$ finite. Then exactly one of the following holds:
  • $\Gamma=\{1\}$, and $(M,g,f)$ is the Bryant soliton.
  • $\Gamma=\mathbb{Z}_{k}$ for some $k\geq 3$, and $(M,g,f)\in\mathcal{A}_{1,k}$.
  • $\Gamma=\tilde{D}_{k}$ for some $k\geq 2$, and $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
We highlight one further consequence of this. Bamler et al. [bamler-3] showed that any tangent flow at infinity of a 4-dimensional steady GRS singularity model is one of
(1.12) $\mathbb{R}^{4}/\Gamma\text{ (but not }\mathbb{R}^{4}\text{)},\quad(\mathbb{S}^{3}/\Gamma)\times\mathbb{R},\quad\mathbb{S}^{2}\times\mathbb{R}^{2},\quad((\mathbb{S}^{2}\times\mathbb{R})/\mathbb{Z}_{2})\times\mathbb{R},$
where $\Gamma$ is a finite subgroup of $O(4)$. Moreover, if one tangent flow at infinity is $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$, then the soliton is asymptotically $\mathbb{S}^{3}/\Gamma$-cylindrical. Hence, Corollary 1.12 classifies all steady GRS singularity models with a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$, where $\Gamma$ is a finite subgroup of $SU(2)$.
1.4. Summary of results
For the reader’s convenience, we gather our classification results into Table 1 below.
$\mathcal{S}$ Asymptotically $\mathcal{S}$-cylindrical steady GRSs Theorem
$\mathbb{S}^{n-1}$ ($n\geq 3$) Bryant soliton 1.4
$\mathbb{RP}^{n-1}$ ($n\geq 3$) None 1.6
$L_{m,k}$ ($m\geq 1$, $k\geq 3$) Appleton solitons in $\mathcal{A}_{m,k}$ 1.8
$SU(2)/\Gamma$ ($\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$) None 1.10
$SU(2)/\tilde{D}_{k}$ ($k\geq 2$) $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.11
Table 1. Classification of asymptotically $\mathcal{S}$-cylindrical steady GRSs for various spherical space forms $\mathcal{S}$.
We also have Corollary 1.5 and Corollary 1.12 as consequences in dimension 4.
1.5. Remarks on the orbifold case
While this paper exclusively works with smooth Riemannian manifolds, a robust singularity theory of Ricci flow in dimensions $\geq 4$ requires objects of lower regularity. Indeed, Bamler’s compactness theory [bamler-a, bamler-b, bamler-c] shows that finite-time singularity models of Ricci flow are orbifold Ricci solitons with singular locus of codimension $\geq 4$.
It is therefore of interest to see if our results generalize to the orbifold setting, and we believe they do. Our proofs use local properties like the maximum principle and the analyticity of Ricci solitons, which generalize directly to orbifolds as one can pass to local uniformizing covers and apply results for smooth manifolds there. We also use the standard formalism for cohomogeneity one manifolds, of which a version for orbifolds exists [gonzalez-alvaro-phd, borzellino-phd]. To this end, the only major change occurs in Proposition 4.2(d), where $K/H$ is now allowed to be a spherical space form $\mathbb{S}^{\ell}/\Gamma$. Naively applying this change to our proofs, it appears that quotients of the Bryant soliton are the only new possible steady GRSs that arise in the expanded category of smooth Riemannian orbifolds. Hence, we conjecture:
Conjecture 1.13.
Table 1 correctly classifies asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth Riemannian orbifolds for certain $\mathcal{S}$, provided that one includes the appropriate Bryant soliton quotient as an extra possibility in each row.
Verifying this conjecture requires checking if global and asymptotic properties of steady GRSs on Riemannian manifolds carry over to orbifolds. For instance, we used connectedness at infinity (Theorem 2.4) whose proof does not generalize directly to orbifolds. Instead, this property is separately established in [bamler-3]*Appendix A for 4-dimensional steady GRS singularity models with a tangent flow at infinity of the form $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$.
In a recent article [deng-orbifold], Deng proved that any asymptotically $\mathcal{S}$-cylindrical steady GRS on a smooth orbifold with positive sectional curvature must be a quotient of the Bryant soliton. Unlike for manifolds, it is impossible to drop the curvature hypothesis (at least entirely), as the solitons in $\mathcal{A}_{m,k}$ would furnish counterexamples. The reason our methods cannot apply here is because Deng employs a reduction to manifolds that fails without curvature conditions. Specifically, Deng obtains a covering of the orbifold GRS $M$ by an asymptotically cylindrical steady GRS on a manifold $\tilde{M}$ with positive sectional curvature. Then Brendle’s uniqueness result [brendle2014rotational] (alternatively Theorem 1.4) applies to $\tilde{M}$. However, the existence of such a cover requires the singular locus of $M$ to consist of a single point, which demands at least the positivity of Ricci curvature (see [deng-orbifold]*§4).
1.6. Outline of the paper
In §2, we collect relevant facts (due to other authors) about asymptotically $\mathcal{S}$-cylindrical steady GRSs. In particular, any such soliton $(M,g,f)$ opens up like a paraboloid, and the level sets of $f$ converge smoothly (after a certain rescaling) to a spherical space form $(\mathcal{S},\bar{g}_{\infty})$.
In §3, we show that each Killing field $\bar{U}$ on $(\mathcal{S},\bar{g}_{\infty})$ induces a Killing field $W$ on $(M,g)$ (Theorem 3.1). Moreover, $W$ restricts to a Killing field on any level set of $f$. Following Brendle [brendle2013rotational, brendle2014rotational], this reduces to proving a Liouville-type theorem for symmetric 2-tensors $h$ solving the elliptic equation
(1.13) $Lh=\Delta h+\nabla_{\nabla f}h+2\operatorname{Rm}(h)=0,$
where $\operatorname{Rm}(h)_{ik}=\operatorname{Rm}_{ijkl}h^{jl}$ is the action of the curvature tensor. Brendle and Zhao–Zhu [zz] carried this out under extra curvature hypotheses. In contrast, our Liouville theorem (Proposition 3.6) is specific to the tensor $h=\mathcal{L}_{W}g$ but requires no curvature hypotheses. The proof of Proposition 3.6 uses parts of Brendle’s original blowdown argument, decay estimates of Dung–Sung [dung-sung] and Zhao–Zhu [zz], tensor identities on Ricci solitons (where $h$ being a Lie derivative is essential), maximum principles, and standard integration-by-parts arguments.
In §4, we use the results of §3 to establish global symmetry principles. Under mild conditions, we show that the isometry group of each level set of $f$ acts isometrically on the whole of $(M,g)$, and this action preserves all level sets of $f$ (Lemma 4.1). Moreover, if the level sets are homogeneous, then the orbit space is one-dimensional (Proposition 4.2). In this case, standard facts about cohomogeneity one manifolds apply to pin down the global structure of $(M,g,f)$.
In §5, we prove the uniqueness of the Bryant soliton (Theorem 1.4) following [brendle2013rotational, brendle2014rotational, zz]. Given an asymptotically cylindrical steady GRS $(M^{n},g,f)$, we use the results of §3 to find $\frac{n(n-1)}{2}$ Killing fields on each far enough level set of $f$, which implies these level sets are isometric to round spheres. From this, we deduce that $(M,g)$ is globally rotationally symmetric. This last step was not detailed by Brendle and Zhao–Zhu; we give its proof for completeness, and also because a slight generalization of it (Proposition 5.2) is needed. We also prove the non-existence of asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRSs (Theorem 1.6). This proceeds almost identically to the proof of Theorem 1.4, except the last step uses Proposition 5.2 to obtain a contradiction.
In §6, we prove the uniqueness of Appleton’s solitons (Theorem 1.8). Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$, $k\geq 3$. The results of §3 induce $(m+1)^{2}$ Killing fields on each far enough level set of $f$. Using the classification of homogeneous metrics on spheres [ziller-homog] and some group theory, we deduce that the level sets are isometric to quotiented Berger spheres $(L_{m,k},g_{a,b})$. The results of §4 imply that $(M,g)$ has a cohomogeneity one isometric action of $U(m+1)$, and is specified outside a submanifold $X$ by Appleton’s ODE system (1.7)–(1.9). The cohomogeneity one formalism classifies the possible diffeomorphism types of $X$; in our case, we find that the metric completes smoothly if and only if $X\cong\mathbb{CP}^{m}$ and the boundary conditions (1.10) are met. We conclude that $(M,g,f)\in\mathcal{A}_{m,k}$.
In §7, we finish the classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs (Theorems 1.10 and 1.11). Here we also describe the solitons $A/\mathbb{Z}_{2}$ (for $A\in\mathcal{A}_{1,2k}$) in more detail. Other than this, the ideas and methods involved are similar to those used in §5 and §6, so we omit an outline for brevity.
Appendix A contains a proof that Appleton’s solitons in $\mathcal{A}_{m,k}$ are asymptotically $L_{m,k}$-cylindrical.
Acknowledgements
The author is grateful to William Minicozzi and Tristan Ozuch for their insights and continual support, as well as Yonghwan Kim for inspiring discussions. The author is supported by a Croucher Scholarship.
2. Geometry of asymptotically $\mathcal{S}$-cylindrical steady GRSs
Throughout this section, $(M^{n},g,f)$ denotes an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}$ is a spherical space form of dimension $n-1$.
2.1. Steady soliton identities
Recall that the steady GRS equation is
(2.1) $\operatorname{Ric}=\mathrm{Hess}_{f}.$
Denote by $R$ the scalar curvature of $(M,g)$. Standard identities for steady GRSs give $R\geq 0$ and $R+|\nabla f|^{2}\equiv c\geq 0$, and if equality holds in either then $\operatorname{Ric}=0$. Clearly, asymptotically $\mathcal{S}$-cylindrical steady GRSs are not Ricci flat, so
(2.2) $R>0$
and by scaling the metric we may arrange that $c=1$; thus
(2.3) $R+|\nabla f|^{2}=1.$
Our convention for the Laplacian on any tensor is $\Delta=-\nabla^{*}\nabla$. The drift Laplacian $\Delta_{f}$ on tensors is defined by
(2.4) $\Delta_{f}=\Delta+\nabla_{\nabla f}.$
This is self-adjoint with respect to the weighted volume form $e^{f}\mathrm{dvol}_{g}$. Under the normalization (2.3), one has the soliton identity
(2.5) $\Delta_{f}f=1$
which will be repeatedly used throughout.
2.2. Geometric properties
Let $r:M\to\mathbb{R}$ be the distance function from a fixed point with respect to $g$.
Lemma 2.1 ([CDN]*§2, also [bamler-3]*Lemma 3.14).
We have $\frac{f(x)}{r(x)}\to 1$ as $r(x)\to\infty$.
In particular, $f$ is proper and bounded from below. As (2.1) and (2.3) are invariant under translating $f$, we will henceforth assume $\min_{M}f=1$.
Lemma 2.2.
For sufficiently large $\rho>0$, we have $\nabla f\neq 0$ on $\{f\geq\rho\}$. Hence, $\{f=\rho\}$ is a smooth hypersurface in $M$.
Proof.
By [deng-zhu]*Lemma 6.5, asymptotic $\mathcal{S}$-cylindricity implies that $R=o(1)$ as $r\to\infty$. Then the claim follows from (2.3) and Lemma 2.1. ∎
Next, we state the roundness estimates on the level sets of $f$ which were first established by Brendle [brendle2014rotational] and subsequently improved by Zhao–Zhu [zz].
Proposition 2.3 ([zz]*Propositions 2.5 and 2.7).
Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal frame such that $e_{n}=\frac{\nabla f}{|\nabla f|}$. There exists $\delta>0$ such that on $M$, we have
(2.6) $fR$ $=\frac{n-1}{2}+\mathcal{O}(r^{-\frac{1}{2}-\delta})$
and
(2.7) $f\operatorname{Rm}_{ijkl}$ $=\frac{1}{2(n-2)}(g_{ik}-\partial_{i}f\partial_{k}f)(g_{jl}-\partial_{j}f\partial_{l}f)$
(2.8) $\quad-\frac{1}{2(n-2)}(g_{il}-\partial_{i}f\partial_{l}f)(g_{jk}-\partial_{j}f\partial_{k}f)+\mathcal{O}(r^{-\frac{1}{2}-\delta}).$
For the full proof of this result, see Propositions 2.1–2.5 in [brendle2014rotational] followed by Section 2.1 of [zz]. Once $f$ and $r$ are known to be comparable (as in Lemma 2.1), these proofs only use asymptotic cylindricity and do not require any curvature hypotheses. They carry over verbatim to asymptotically $\mathcal{S}$-cylindrical steady GRSs.
It is clear that $(M,g)$ does not split off a line. Hence:
Theorem 2.4 ([mw-smms]*Corollary 1.1).
$M$ has exactly one end.
Corollary 2.5.
The hypersurfaces $\{f=\rho\}$ are diffeomorphic to $\mathcal{S}$ for all large $\rho$.
Proof.
By Lemma 2.1 and Lemma 2.2, the level sets $\{f=\rho\}$ are smooth hypersurfaces bounding the end(s) of $M$. Then by Theorem 2.4, $\{f=\rho\}$ is connected for all large $\rho$.
Take a sequence of points $p_{m}$ going to infinity, and let $r_{m}=f(p_{m})$, $\hat{g}^{(m)}(t)=r_{m}^{-1}\Phi_{r_{m}t}^{*}g$. By [deng-zhu]*Lemma 6.5, the Ricci curvature of the level sets of $f$ decays linearly in $r$. Then by the Bonnet–Myers theorem, the $g$-diameter of $\{f=r_{m}\}$ is bounded by $C\sqrt{r_{m}}$, so the $\hat{g}^{(m)}(t)$-diameter of $\{f=r_{m}\}$ is uniformly bounded over all $m$ and over all $t$ in a compact interval within $(0,1)$.
By asymptotic $\mathcal{S}$-cylindricity, the sequence $(M,\hat{g}^{(m)}(t),p_{m})$ converges in the Cheeger–Gromov sense to a shrinking $\mathcal{S}$-cylinder (1.4) for $t\in(0,1)$. Moreover, the vector fields $\sqrt{r_{m}}\nabla f$ converge to the axial vector field $\frac{\partial}{\partial z}$ on $\mathcal{S}\times\mathbb{R}$. Using this, and the assertions on the $\hat{g}^{(m)}(t)$-diameter of $\{f=r_{m}\}$ from the last paragraph, it is easy to see that $\{f=r_{m}\}$ is diffeomorphic to $\mathcal{S}$ for all large $m$. ∎
By the above results, $f$ is a coordinate on the end of $M$ with level sets diffeomorphic to $\mathcal{S}$. Let $F_{\rho}:\mathcal{S}\to\{f=\rho\}\subset M$ be a one-parameter family of diffeomorphisms such that $\frac{\partial}{\partial\rho}F_{\rho}=\frac{\nabla f}{|\nabla f|^{2}}$. Define Riemannian metrics $g_{\rho}$ on $\mathcal{S}$ by $g_{\rho}=F_{\rho}^{*}g$. Then the end of $(M,g)$ is isometric to
(2.9) $\left(I\times\mathcal{S},\frac{df^{2}}{|\nabla f|^{2}}+g_{f}\right),$
where $I$ is an interval of the form $[\rho_{0},\infty)$. Following [zz]*§2.2, we have the following convergence and closeness bounds:
Lemma 2.6.
There exists $\delta>0$ such that for each $\ell\geq 0$, we have
(2.10) $\left\lVert\frac{1}{2(n-2)\rho}g_{\rho}-\bar{g}_{\infty}\right\rVert_{C^{\ell}(\mathcal{S},\bar{g}_{\infty})}\leq\mathcal{O}(\rho^{-\frac{1}{2}-\frac{1}{2}\delta})$
as $\rho\to\infty$, where $\bar{g}_{\infty}$ is a metric on $\mathcal{S}$ with constant sectional curvature 1.
Thus, $(M,g)$ opens up like a paraboloid, and has volume growth of order $r^{\frac{n+1}{2}}$.
In what follows, given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, the metric $\bar{g}_{\infty}$ is understood to come from Lemma 2.6. We will also frequently identify level sets $\{f=\rho\}$ with $(\mathcal{S},g_{\rho})$ and the end of $(M,g)$ with (2.9).
3. Killing fields
Given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, we will show in this section that every symmetry of the asymptotic cross-section $(\mathcal{S},\bar{g}_{\infty})$ manifests as a genuine symmetry of $(M,g)$. More precisely, we prove:
Theorem 3.1.
Let $(M^{n},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS. Let $\bar{U}$ be a Killing field on $(\mathcal{S},\bar{g}_{\infty})$, and $U$ a smooth vector field on $M$ whose restriction to every sufficiently far level set $\{f=\rho\}\cong\mathcal{S}$ satisfies
(3.1) $U|_{\{f=\rho\}}=\bar{U}.$
Then there exists a vector field $W$ on $(M,g)$ such that
  1. (a)
    $|W-U|\leq\mathcal{O}(1)$.
  2. (b)
    $[W,\nabla f]=0$.
  3. (c)
    $W$ is a Killing field for $(M,g)$, i.e. $\mathcal{L}_{W}g=0$.
  4. (d)
    $W$ is tangent to the level sets of $f$, i.e. $\langle W,\nabla f\rangle=0$. Hence, for all sufficiently large $\rho>0$, $W$ is a Killing field for the level set $(\mathcal{S},g_{\rho})$.
Applying this multiple times, we obtain the following corollary.
Corollary 3.2.
Let $(M^{n},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS. Let $\{\bar{U}_{i}\}_{i=1}^{d}$ be a basis for the space of Killing fields on $(\mathcal{S},\bar{g}_{\infty})$. Then there exists a set of vector fields $\{W_{i}\}_{i=1}^{d}$ on $M$ such that
  1. (a)
    $[W_{i},\nabla f]=0$.
  2. (b)
    $\langle W_{i},\nabla f\rangle=0$.
  3. (c)
    $W_{i}$ is a Killing field for $(M,g)$ and for each sufficiently far level set of $f$.
  4. (d)
    For all large $\rho>0$, $\{W_{i}\}_{i=1}^{d}$ is an $\mathbb{R}$-linearly independent set of vector fields on the hypersurface $\{f=\rho\}$. Moreover, if $\{\bar{U}_{i}\}_{i=1}^{d}$ spans the tangent space to $\mathcal{S}$ at every point, then $\{W_{i}\}_{i=1}^{d}$ spans the tangent space to $\{f=\rho\}$ at every point for all large $\rho$.
Proof.
Applying Theorem 3.1 to each $\bar{U}_{i}$ gives vector fields $\{W_{i}\}_{i=1}^{d}$ satisfying (a), (b), and (c). Moreover, for each $i$,
(3.2) $|W_{i}-U_{i}|\leq\mathcal{O}(1),$
where $U_{i}$ is a vector field on $M$ such that $U_{i}|_{\{f=\rho\}}=\bar{U}_{i}$ for all large $\rho>0$. Using (3.2) and Lemma 2.6, and identifying $\{f=\rho\}\cong\mathcal{S}$, we see that
(3.3) $\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}$ $\leq\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}-\frac{1}{2(n-2)\rho}g_{\rho}}^{2}+\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\frac{1}{2(n-2)\rho}g_{\rho}}^{2}$
(3.4) $\leq\mathcal{O}(\rho^{-\frac{1}{2}-\frac{1}{2}\delta})\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}+\mathcal{O}(\rho^{-1}).$
Hence,
(3.5) $\sup_{\{f=\rho\}}|W_{i}-\bar{U}_{i}|_{\bar{g}_{\infty}}^{2}=\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}\leq\mathcal{O}(\rho^{-1}).$
Since $\{\bar{U}_{i}\}_{i=1}^{d}$ is $\mathbb{R}$-linearly independent on $\mathcal{S}\cong\{f=\rho\}$, and this is an open condition with respect to the $C^{0}(\bar{g}_{\infty})$ topology, it follows that $\{W_{i}\}_{i=1}^{d}$ is $\mathbb{R}$-linearly independent on $\{f=\rho\}$ for all large $\rho$. Likewise, if $\{\bar{U}_{i}\}_{i=1}^{d}$ spans the tangent space to $\mathcal{S}\cong\{f=\rho\}$ at every point, then $\{W_{i}\}_{i=1}^{d}$ spans the tangent space to $\{f=\rho\}$ at every point, whenever $\rho$ is large enough. ∎
3.1. Reduction to a Liouville theorem for symmetric 2-tensors
For the rest of this section, we fix an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$. The goal is to prove Theorem 3.1. First, we pull back Killing fields from $(\mathcal{S},\bar{g}_{\infty})$ to obtain approximate Killing fields on $(M,g)$.
Proposition 3.3 ([zz]*Proposition 3.1).
Let $\bar{U}$ be a Killing field on $(\mathcal{S},\bar{g}_{\infty})$. Let $U$ be a smooth vector field on $M$ whose restriction to every sufficiently far level set $\{f=\rho\}$ satisfies
(3.6) $U|_{\{f=\rho\}}=\bar{U}.$
Then for some $\delta>0$, we have
(3.7) $|\mathcal{L}_{U}g|$ $\leq\mathcal{O}(r^{-\frac{1}{2}-\frac{1}{2}\delta}),$
(3.8) $|\Delta_{f}U|$ $\leq\mathcal{O}(r^{-1-\frac{1}{4}\delta}).$
Note that $|U|\leq\mathcal{O}(r^{\frac{1}{2}})$ by Lemma 2.6. Correcting $U$ by a bounded vector field yields a candidate Killing field on $(M,g)$:
Proposition 3.4.
Let $U$ be as in Proposition 3.3. Then there is a smooth vector field $W$ on $M$ such that
(3.9) $\Delta_{f}W$ $=0,$
(3.10) $|W-U|$ $\leq\mathcal{O}(1),$
(3.11) $|W|$ $\leq\mathcal{O}(r^{\frac{1}{2}}),$
(3.12) $|\mathcal{L}_{W}g|$ $\leq\mathcal{O}(r^{-\frac{1}{2}}).$
Proof.
From Proposition 3.3, we have $|\Delta_{f}U|\leq\mathcal{O}(r^{-1-\frac{1}{4}\delta})$. Following the proofs of Lemma 3.2 and Theorem 3.3 in [zz], there exists a solution $V$ of
(3.13) $\Delta_{f}V=-\Delta_{f}U$
such that $|V|\leq\mathcal{O}(1)$ and $|\nabla V|\leq\mathcal{O}(r^{-\frac{1}{2}})$. Let $W=U+V$. Using the bounds on $|V|$ and $|\nabla V|$ together with Proposition 3.3, the assertions follow. ∎
Define the operator $L$ on symmetric 2-tensors $h$ by
(3.14) $Lh=\Delta_{f}h+2\operatorname{Rm}(h),$
where $\operatorname{Rm}(h)_{ik}=\operatorname{Rm}_{ijkl}h^{jl}$. As $\operatorname{Ric}=\mathrm{Hess}_{f}$, one easily checks that
(3.15) $Lh=\Delta_{L}h+\mathcal{L}_{\nabla f}h,$
where $\Delta_{L}$ is the Lichnerowicz Laplacian defined by
(3.16) $\Delta_{L}h_{ik}=\Delta h_{ik}+2\operatorname{Rm}_{ijkl}h^{jl}-\operatorname{Ric}_{i}^{l}h_{kl}-\operatorname{Ric}_{k}^{l}h_{il}.$
Brendle made the following crucial observations:
Theorem 3.5 ([brendle2013rotational]*§4).
Let $W$ be a vector field satisfying $\Delta_{f}W=0$. Then $h=\mathcal{L}_{W}g$ satisfies $Lh=0$. In particular, taking $W=\nabla f$ gives $L\operatorname{Ric}=0$.
Brendle [brendle2013rotational, brendle2014rotational] and Zhao–Zhu [zz] proved Liouville theorems for solutions of $Lh=0$ under certain curvature hypotheses. In contrast, our Liouville theorem is specific to $h=\mathcal{L}_{W}g$, where $W$ is from Proposition 3.4, and requires no curvature assumptions:
Proposition 3.6.
Let $W$ be a vector field on $M$ such that
(3.17) $\Delta_{f}W$ $=0,$
(3.18) $|W|$ $\leq\mathcal{O}(r^{\frac{1}{2}}),$
(3.19) $|\mathcal{L}_{W}g|$ $\leq\mathcal{O}(r^{-\frac{1}{2}}).$
Then $\mathcal{L}_{W}g=\lambda\operatorname{Ric}$ for some $\lambda\in\mathbb{R}$.
Following the proof of [brendle2013rotational]*Theorem 7.1, we show how this implies Theorem 3.1:
Proof of Theorem 3.1.
Let $\bar{U}$ and $U$ be as in the statement of the theorem. By Propositions 3.4 and 3.6, there exists a vector field $W$ on $M$ such that $|W-U|\leq\mathcal{O}(1)$ and $\mathcal{L}_{W}g=\lambda\operatorname{Ric}$. Let
(3.20) $\tilde{W}=W-\frac{\lambda}{2}\nabla f.$
We claim that $\tilde{W}$ satisfies properties (a)–(d) in the theorem. Since $\operatorname{Ric}=\mathrm{Hess}_{f}=\frac{1}{2}\mathcal{L}_{\nabla f}g$, it follows that $\mathcal{L}_{\tilde{W}}g=0$. Since $\nabla f$ is bounded, we have $|\tilde{W}-U|\leq\mathcal{O}(1)$. This proves (a) and (c).
By (2.5) and the weighted Bochner formula, we have $\Delta_{f}\nabla f=0$, so $\Delta_{f}\tilde{W}=0$. As in [brendle2013rotational, brendle2014rotational], the relation $\mathcal{L}_{\tilde{W}}g=0$ implies $\Delta\tilde{W}+\operatorname{Ric}(\tilde{W})=0$. Thus, $[\tilde{W},\nabla f]=\operatorname{Ric}(\tilde{W})-\nabla_{\nabla f}\tilde{W}=-\Delta_{f}\tilde{W}=0$, establishing property (b). Finally, we have
(3.21) $\nabla^{2}(\mathcal{L}_{\tilde{W}}f)$ $=\mathcal{L}_{\tilde{W}}(\nabla^{2}f)=\frac{1}{2}\mathcal{L}_{\tilde{W}}(\mathcal{L}_{\nabla f}g)=\frac{1}{2}\mathcal{L}_{\nabla f}(\mathcal{L}_{\tilde{W}}g)=0,$
so the function $\mathcal{L}_{\tilde{W}}f=\langle\tilde{W},\nabla f\rangle$ is constant. Since $f$ attains a minimum by Lemma 2.1, we conclude that $\langle\tilde{W},\nabla f\rangle$ vanishes identically. From this, and the identification (2.9) of the end of $(M,g)$, one checks (e.g. by a computation in coordinates) that $\mathcal{L}_{\tilde{W}}g_{\rho}=\mathcal{L}_{\tilde{W}}g=0$ for each large $\rho$. Thus $\tilde{W}$ is a Killing field on $(\mathcal{S},g_{\rho})$. ∎
3.2. Superpolynomial decay
The remainder of this section is spent proving Proposition 3.6. We begin with two computations regarding solutions of $Lh=0$.
Lemma 3.7.
Let $h$ be a symmetric 2-tensor on $M$ satisfying $Lh=0$.
  1. (a)
    For all sufficiently large $\ell>0$, there exists a compact set $K_{\ell}\subset M$ independent of $h$ such that
    (3.22) $\Delta_{f}(f^{\ell}|h|^{2})\geq 0\quad\text{on }M\setminus K_{\ell}.$
  2. (b)
    For all $\varepsilon\in(0,1)$, there exists a compact set $K_{\varepsilon}\subset M$ independent of $h$ such that
    (3.23) $\Delta_{f}(e^{\varepsilon f}|h|^{2})\geq 0\quad\text{on }M\setminus K_{\varepsilon}.$
Proof.
Since $Lh=0$, we have
(3.24) $\Delta_{f}|h|^{2}$ $=2\langle\Delta_{f}h,h\rangle+2|\nabla h|^{2}=-4\langle\operatorname{Rm}(h),h\rangle+2|\nabla h|^{2}.$
Let $\ell>0$. Using this and the soliton identity $\Delta_{f}f=1$, we get
(3.25) $\Delta_{f}(f^{\ell}|h|^{2})$ $=(\Delta_{f}f^{\ell})|h|^{2}+f^{\ell}\Delta_{f}|h|^{2}+2\langle\nabla f^{\ell},\nabla|h|^{2}\rangle$
(3.26) $=\left(\ell f^{\ell-1}+\ell(\ell-1)f^{\ell-2}|\nabla f|^{2}\right)|h|^{2}$
(3.27) $\quad+f^{\ell}\left(2|\nabla h|^{2}-4\langle\operatorname{Rm}(h),h\rangle\right)+2\ell f^{\ell-1}\langle\nabla f,\nabla|h|^{2}\rangle$
(3.28) $\geq\ell f^{\ell-1}|h|^{2}+\ell(\ell-1)f^{\ell-2}|\nabla f|^{2}|h|^{2}+2f^{\ell}|\nabla h|^{2}$
(3.29) $\quad-4f^{\ell}|{\operatorname{Rm}}||h|^{2}-4\ell f^{\ell-1}|\nabla f||h||\nabla|h||.$
By Proposition 2.3, we have
(3.30) $|{\operatorname{Rm}}|\leq Cf^{-1}.$
Moreover, by the absorbing inequality and Kato’s inequality, we have
(3.31) $2|\nabla f||h||\nabla|h||\leq\frac{f}{\ell}|\nabla h|^{2}+\frac{\ell}{f}|\nabla f|^{2}|h|^{2}.$
Using these facts in (3.29) gives
(3.32) $\Delta_{f}(f^{\ell}|h|^{2})$ $\geq(\ell-C)f^{\ell-1}|h|^{2}-(\ell+\ell^{2})f^{\ell-2}|\nabla f|^{2}|h|^{2}$
(3.33) $=f^{\ell-2}|h|^{2}\left[(\ell-C)f-(\ell+\ell^{2})|\nabla f|^{2}\right].$
Since $|\nabla f|^{2}\leq 1$, and $f\to+\infty$ at infinity, it follows that if $\ell>C$, the right-hand side is nonnegative outside a compact set of $M$. This proves (a).
The proof of (b) is similar. For any $\varepsilon\in(0,1)$, we compute
(3.34) $\Delta_{f}(e^{\varepsilon f}|h|^{2})$ $=\left(\varepsilon e^{\varepsilon f}+\varepsilon^{2}e^{\varepsilon f}|\nabla f|^{2}\right)|h|^{2}$
(3.35) $\quad+e^{\varepsilon f}\left(2|\nabla h|^{2}-4\langle\operatorname{Rm}(h),h\rangle\right)+2\varepsilon e^{\varepsilon f}\langle\nabla f,\nabla|h|^{2}\rangle$
(3.36) $\geq e^{\varepsilon f}(\varepsilon+\varepsilon^{2}|\nabla f|^{2})|h|^{2}+2e^{\varepsilon f}|\nabla h|^{2}$
(3.37) $\quad-4e^{\varepsilon f}|{\operatorname{Rm}}||h|^{2}-4\varepsilon e^{\varepsilon f}|\nabla f||h||\nabla|h||.$
By the absorbing inequality and Kato’s inequality, we have
(3.38) $2|\nabla f||h||\nabla|h||$ $\leq\varepsilon^{-1}|\nabla h|^{2}+\varepsilon|\nabla f|^{2}|h|^{2}.$
Using this in (3.37), we obtain
(3.39) $\Delta_{f}(e^{\varepsilon f}|h|^{2})$ $\geq e^{\varepsilon f}(\varepsilon-\varepsilon^{2}|\nabla f|^{2})|h|^{2}-4e^{\varepsilon f}|{\operatorname{Rm}}||h|^{2}$
(3.40) $=e^{\varepsilon f}|h|^{2}(\varepsilon-\varepsilon^{2}-4|{\operatorname{Rm}}|).$
Since $\varepsilon\in(0,1)$ and $|{\operatorname{Rm}}|$ decays by (3.30), the right-hand side is nonnegative outside a compact set of $M$. This proves (b). ∎
The next two results are due to Brendle and Zhao–Zhu, respectively.
Proposition 3.8 ([brendle2014rotational]*Lemma 4.1).
Consider the shrinking cylinder $(\mathcal{S}\times\mathbb{R},\bar{g}(t))$, $t\in(0,1)$, where $\bar{g}(t)$ is given by (1.4). Let $\bar{h}(t)$, $t\in(0,1)$, be a one-parameter family of symmetric 2-tensors which solve the parabolic equation
(3.41) $\frac{\partial}{\partial t}\bar{h}(t)=\Delta_{L,\bar{g}(t)}\bar{h}(t),$
where $\Delta_{L,\bar{g}(t)}$ is the Lichnerowicz Laplacian (see (3.16)). Moreover, suppose $\bar{h}(t)$ is invariant under translations along the axis of the cylinder, and
(3.42) $|\bar{h}(t)|_{\bar{g}(t)}\leq(1-t)^{-\ell}$
for all $t\in(0,\frac{1}{2}]$ and some $\ell>0$. Then there exists $N>0$ depending on $\ell$ such that
(3.43) $\inf_{\lambda\in\mathbb{R}}\sup_{\mathcal{S}\times\mathbb{R}}|\bar{h}(t)-{\lambda\operatorname{Ric}_{\bar{g}(t)}}|_{\bar{g}(t)}\leq N(1-t)^{\frac{1}{2(n-2)}-\frac{1}{2}}$
for all $t\in[\frac{1}{2},1)$.
Proof.
The proof of [brendle2014rotational]*Lemma 4.1 applies verbatim after replacing every occurrence of $(\mathbb{S}^{n-1},g_{\mathbb{S}^{n-1}})$ there with $(\mathcal{S},\bar{g}_{\mathcal{S}})$. (The proof uses spectral properties of the Laplacian which still hold on $(\mathcal{S},\bar{g}_{\mathcal{S}})$.) ∎
Proposition 3.9 ([zz]*Theorem 4.3).
If $Lh=0$ on $M$ and $|h|=\mathcal{O}(r^{-\frac{1}{2}})$, then there exists $\lambda\in\mathbb{R}$ such that $|h-{\lambda\operatorname{Ric}}|=\mathcal{O}(r^{-\ell})$ for every $\ell>0$.
Proof sketch.
Due to differences in organization, we will outline the proof, omitting steps which can essentially be copied down from the original source.
Let $h$ satisfy $Lh=0$ and $|h|=\mathcal{O}(r^{-\frac{1}{2}})$. Consider the function
(3.44) $A(\rho)=\inf_{\lambda\in\mathbb{R}}\sup_{\{f=\rho\}}|h-{\lambda\operatorname{Ric}}|.$
Let $r_{m}\to\infty$ be a sequence of numbers, and denote by $\lambda_{m}$ the minimizing $\lambda$’s associated with $A(r_{m})$.
By Theorem 3.5, we have $L(h-\lambda_{m}\operatorname{Ric})=0$ for each $m$. Then by Lemma 3.7 and the maximum principle, for each large $\ell>0$ there exists $\rho_{\ell}>0$ such that for all $m$,
(3.45) $\sup_{\{\rho_{\ell}\leq f\leq r_{m}\}}f^{\ell}|h-{\lambda_{m}\operatorname{Ric}}|^{2}\leq\max\left\{\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda_{m}\operatorname{Ric}}|^{2},r_{m}^{\ell}\sup_{\{f=r_{m}\}}|h-{\lambda_{m}\operatorname{Ric}}|^{2}\right\}.$
Taking square roots and redefining $\rho_{\ell}$ appropriately, this becomes
(3.46) $\sup_{\{\rho_{\ell}\leq f\leq r_{m}\}}f^{\ell}|h-\lambda_{m}\operatorname{Ric}|\leq\max\left\{\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda_{m}\operatorname{Ric}}|,r_{m}^{\ell}\sup_{\{f=r_{m}\}}|h-{\lambda_{m}\operatorname{Ric}}|\right\}.$
We divide into two cases:
Case 1: there exists $\ell>0$ and a subsequence of $r_{m}$ (still denoted $r_{m}$) such that
(3.47) $\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda_{m}\operatorname{Ric}}|\leq r_{m}^{\ell}\sup_{\{f=r_{m}\}}|h-{\lambda_{m}\operatorname{Ric}}|.$
This is Case 1 in the proof of [zz]*Theorem 4.3. Following the argument there (which uses Proposition 3.8), we see that the $\lambda_{m}$’s are all equal to a constant $\lambda$, and
(3.48) $h=\lambda\operatorname{Ric}\quad\text{on }\{f\geq\rho_{\ell}\}.$
This clearly implies $|h-{\lambda\operatorname{Ric}}|\leq\mathcal{O}(r^{-\ell})$ for every $\ell>0$.
Case 2: for all $\ell>0$, there exists $R_{\ell}>0$ such that for all $r_{m}>R_{\ell}$,
(3.49) $\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda_{m}\operatorname{Ric}}|>r_{m}^{\ell}\sup_{\{f=r_{m}\}}|h-{\lambda_{m}\operatorname{Ric}}|.$
Then by (3.46), for each $\ell>0$, and for all $m$ such that $r_{m}>R_{\ell}$, one has
(3.50) $\sup_{\{\rho_{\ell}\leq f\leq r_{m}\}}f^{\ell}|h-{\lambda_{m}\operatorname{Ric}}|\leq\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda_{m}\operatorname{Ric}}|.$
Following the arguments in Case 2 in the proof of [zz]*Theorem 4.3 (in particular Claim 4.4 there), we conclude that the sequence $\lambda_{m}$ is uniformly bounded, so a subsequence of the $\lambda_{m}$ converges to some $\lambda\in\mathbb{R}$. Fixing an arbitrary $\ell>0$ in (3.50) and taking $m\to\infty$, we get
(3.51) $\sup_{\{f\geq\rho_{\ell}\}}f^{\ell}|h-{\lambda\operatorname{Ric}}|\leq\rho_{\ell}^{\ell}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda\operatorname{Ric}}|.$
Hence, for all $\rho\geq\rho_{\ell}$,
(3.52) $\sup_{\{f=\rho\}}|h-{\lambda\operatorname{Ric}}|\leq\frac{\rho_{\ell}^{\ell}}{\rho^{\ell}}\sup_{\{f=\rho_{\ell}\}}|h-{\lambda\operatorname{Ric}}|\leq C_{\ell}\rho^{-\ell}$
where $C_{\ell}>0$ depends on $\ell$. So $|h-{\lambda\operatorname{Ric}}|\leq\mathcal{O}(r^{-\ell})$ for each $\ell>0$. ∎
3.3. Superexponential decay
Next, we promote the decay obtained from Proposition 3.9 to superexponential. This will be done using two decay estimates for functions which are drift-subharmonic outside a compact set. The first is a simple application of soliton identities and the maximum principle.
Lemma 3.10.
Suppose $u$ is a function such that $\Delta_{f}u\geq 0$ outside a compact set and $u\to 0$ at infinity. Then $u\leq Ce^{-f}$ on $M$ for some $C>0$.
Proof.
Recall that $R\geq 0$ on any steady GRS. By this and (2.3),
(3.53) $\Delta_{f}e^{-f}=-e^{-f}\Delta_{f}f+e^{-f}|\nabla f|^{2}=e^{-f}(-1+|\nabla f|^{2})=-Re^{-f}\leq 0.$
Choose $\rho>0$ large so that $\Delta_{f}u\geq 0$ on $\{f>\rho\}$. Choose $C>0$ large so that $u-Ce^{-f}<0$ on $\{f=\rho\}$. We have
(3.54) $\Delta_{f}(u-Ce^{-f})\geq 0\quad\text{on }\{f>\rho\},$
and $u-Ce^{-f}\to 0$ at infinity. It follows by the maximum principle that $u-Ce^{-f}\leq 0$ on $\{f>\rho\}$. Suitably increasing $C$, the lemma follows. ∎
By a result of Munteanu and Wang [mw-smms]*Proposition 2.1, any steady GRS supports a weighted Poincaré inequality. In fact,
(3.55) $\frac{1}{4}\int_{M}\phi^{2}e^{f}\leq\int_{M}|\nabla\phi|^{2}e^{f}\quad\text{for all }\phi\in C^{\infty}_{c}(M).$
Also, $M$ has exactly one end by Theorem 2.4. Using these two facts in [dung-sung]*Theorem 2.1, we obtain the following decay estimate.
Lemma 3.11.
Let $u$ be a nonnegative function satisfying $\Delta_{f}u\geq 0$ outside a compact set. Let $r$ be the distance function from a fixed point, and denote by $B_{\rho}$ (for $\rho>0$) the corresponding metric balls. If $u$ satisfies the growth bound
(3.56) $\int_{B_{\rho}}u^{2}e^{-2r}e^{f}=o(\rho)$
as $\rho\to\infty$, then it must satisfy the decay estimate
(3.57) $\int_{B_{\rho+1}\setminus B_{\rho}}u^{2}e^{f}\leq Ce^{-2\rho}$
for some constant $C>0$.
Combining the above results, we have:
Proposition 3.12.
If $Lh=0$ on $M$ and $|h|=\mathcal{O}(r^{-\frac{1}{2}})$, then there exists $\lambda\in\mathbb{R}$ such that $\tilde{h}=h-\lambda\operatorname{Ric}$ satisfies
(3.58) $|\tilde{h}|\leq\mathcal{O}(e^{-Af})\quad\text{for all }A<1,$
and
(3.59) $\int_{\{R<f<2R\}}|\tilde{h}|^{4}e^{4f}\leq Ce^{-\frac{R}{4}}$
for all large $R$, where $C>0$ is independent of $R$.
Proof.
By Proposition 3.9, we can find $\lambda\in\mathbb{R}$ such that $\tilde{h}=h-\lambda\operatorname{Ric}$ satisfies $f^{\ell}|\tilde{h}|^{2}\to 0$ at infinity for all $\ell>0$. Moreover, $L\tilde{h}=0$ by Theorem 3.5. By Lemma 3.7(a), it holds for each sufficiently large $\ell>0$ that
(3.60) $\Delta_{f}(f^{\ell}|\tilde{h}|^{2})\geq 0\quad\text{outside a compact set}.$
Then Lemma 3.10 gives $f^{\ell}|\tilde{h}|^{2}\leq\mathcal{O}(e^{-f})$ for all $\ell>0$. It follows that for any $\varepsilon\in(0,1)$, we have $e^{\varepsilon f}|\tilde{h}|^{2}\to 0$ at infinity. Moreover, by Lemma 3.7(b),
(3.61) $\Delta_{f}(e^{\varepsilon f}|\tilde{h}|^{2})\geq 0\quad\text{outside a compact set}.$
Then Lemma 3.10 gives $e^{\varepsilon f}|\tilde{h}|^{2}\leq\mathcal{O}(e^{-f})$ for all $\epsilon\in(0,1)$. This establishes (3.58).
To prove (3.59), let $u=e^{\frac{3}{4}f}|\tilde{h}|^{2}$. By (3.61), we have
(3.62) $\Delta_{f}u\geq 0\quad\text{outside a compact set}.$
Also, (3.58) implies $|\tilde{h}|^{4}\leq\mathcal{O}(e^{-\frac{7}{2}f})$, so for all metric balls $B_{\rho}$ with $\rho>0$ we have
(3.63) $\int_{B_{\rho}}u^{2}e^{-2r}e^{f}$ $\leq\int_{M}u^{2}e^{f}=\int_{M}|\tilde{h}|^{4}e^{\frac{5}{2}f}\leq C\int_{M}e^{-f}\leq C<\infty,$
because $M$ has volume growth of order $r^{\frac{n+1}{2}}$. By (3.62) and (3.63), we can apply Lemma 3.11 to get
(3.64) $\int_{B_{\rho+1}\setminus B_{\rho}}|\tilde{h}|^{4}e^{\frac{5}{2}f}\leq Ce^{-2\rho}\quad\text{for all sufficiently large }\rho.$
Since $\frac{f}{r}\to 1$ at infinity (Lemma 2.1), we have $\{R<f<R+1\}\subset B_{\frac{9}{8}R}\setminus B_{\frac{7}{8}R}$ for all large $R$. Using (3.64), we estimate
(3.65) $\int_{\{R<f<R+1\}}|\tilde{h}|^{4}e^{\frac{5}{2}f}$ $\leq\int_{(B_{\frac{7}{8}R+1}\setminus B_{\frac{7}{8}R})}|\tilde{h}|^{4}e^{\frac{5}{2}f}+\cdots+\int_{(B_{\frac{9}{8}R}\setminus B_{\frac{9}{8}R-1})}|\tilde{h}|^{4}e^{\frac{5}{2}f}$
(3.66) $\leq Ce^{-2(\frac{7}{8}R)}+Ce^{-2(\frac{7}{8}R+1)}+\cdots+Ce^{-2(\frac{9}{8}R-1)}$
(3.67) $\leq Ce^{-\frac{7}{4}R}.$
Hence,
(3.68) $\int_{\{R<f<2R\}}|\tilde{h}|^{4}e^{4f}$ $\leq\sum_{i=0}^{R-1}\int_{\{R+i<f<R+i+1\}}|\tilde{h}|^{4}e^{\frac{5}{2}f}e^{\frac{3}{2}f}$
(3.69) $\leq C\sum_{i=0}^{R-1}e^{\frac{3}{2}(R+i+1)}e^{-\frac{7}{4}(R+i)}$
(3.70) $\leq Ce^{-\frac{R}{4}}.$
This proves (3.59). ∎
3.4. Vanishing and completion of the proof
For a symmetric 2-tensor $h$, the 1-form $\operatorname{div}_{f}h$ is defined as
(3.71) $\operatorname{div}_{f}h$ $=e^{-f}\operatorname{div}(e^{f}h)=\operatorname{div}h+h(\nabla f,\cdot)$
where $(\operatorname{div}h)_{j}=\nabla^{i}h_{ij}$ is the unweighted divergence. The $L^{2}(e^{f})$-adjoint of $\operatorname{div}_{f}$ is the operator $\operatorname{div}_{f}^{*}$ on vector fields, given by
(3.72) $(\operatorname{div}_{f}^{*}Y)_{ij}=-\frac{1}{2}(\nabla_{i}Y_{j}+\nabla_{j}Y_{i}).$
That is, if $\int_{M}(|Y|^{2}+|\nabla Y|^{2}+|h|^{2}+|\nabla h|^{2})e^{f}<\infty$, then $\int_{M}\langle\operatorname{div}_{f}h,Y\rangle e^{f}=\int_{M}\langle h,\operatorname{div}_{f}^{*}Y\rangle e^{f}$. Note that $\operatorname{div}_{f}^{*}Y=-\frac{1}{2}\mathcal{L}_{Y}g$. We will also need an identity proved in [cm-sing]*Theorem 1.32: for any symmetric 2-tensor $h$ on a steady GRS,
(3.73) $\Delta_{f}\operatorname{div}_{f}h=\operatorname{div}_{f}Lh.$
We now close off this section by proving Proposition 3.6.
Proof of Proposition 3.6.
Let $W$ be a vector field on $M$ such that $\Delta_{f}W=0$, $|W|\leq\mathcal{O}(r^{\frac{1}{2}})$, and $|\mathcal{L}_{W}g|\leq\mathcal{O}(r^{-\frac{1}{2}})$. Write $h=\mathcal{L}_{W}g$. We have $Lh=0$ by Theorem 3.5.
By Proposition 3.12, we can find $\lambda\in\mathbb{R}$ so that $\tilde{h}=h-\lambda\operatorname{Ric}$ satisfies
(3.74) $|\tilde{h}|\leq\mathcal{O}(e^{-Af})\quad\text{for all }A<1$
and
(3.75) $\int_{\{R<f<2R\}}|\tilde{h}|^{4}e^{4f}$ $\leq Ce^{-\frac{R}{4}}$
for sufficiently large $R$, where $C>0$ is independent of $R$. Write $\tilde{W}=W-\frac{\lambda}{2}\nabla f$, so that $|\tilde{W}|\leq\mathcal{O}(r^{\frac{1}{2}})$ and $\tilde{h}=\mathcal{L}_{\tilde{W}}g$. We will show that $\tilde{h}=0$.
By Theorem 3.5, we have $L\tilde{h}=0$. It follows from (3.73) that $\Delta_{f}\operatorname{div}_{f}\tilde{h}=0$.
We claim that $\operatorname{div}_{f}\tilde{h}=0$. This is proved by a standard cutoff argument. For each $R\gg 1$, let $\eta_{R}$ be a smooth function on $M$ such that
(3.76) $\begin{cases}\eta_{R}=1&\text{on }\{f\leq R\},\\ \eta_{R}=0&\text{on }\{f\geq 2R\},\\ 0\leq\eta_{R}\leq 1&\text{on }M,\\ |\nabla\eta_{R}|\leq\frac{C}{R}&\text{on }M,\end{cases}$
where $C>0$ is independent of $R$. Then writing $Z=\operatorname{div}_{f}\tilde{h}$, we have
(3.77) 00 0 $=-\int_{M}\langle\Delta_{f}Z,\eta_{R}^{2}Z\rangle e^{f}=\int_{M}\langle\nabla Z,\nabla(\eta_{R}^{2}Z)\rangle e^{f}$
(3.78) $\geq\int_{M}\eta_{R}^{2}|\nabla Z|^{2}e^{f}-2\int_{M}\eta_{R}|\nabla\eta_{R}||Z||\nabla Z|e^{f}.$
Using the absorbing inequality $2ab\leq\frac{1}{2}a^{2}+2b^{2}$, and rearranging, we get
(3.79) $\frac{1}{2}\int_{M}\eta_{R}^{2}|\nabla Z|^{2}e^{f}\leq 2\int_{M}|\nabla\eta_{R}|^{2}|Z|^{2}e^{f}\leq\frac{C}{R^{2}}\int_{M}|Z|^{2}e^{f}.$
Using the pointwise bound (3.74) together with the gradient estimates of Deruelle [deruelle-asymptotic]*Lemma 2.8 (see also Remark 2.7 there), we get
(3.80) $|Z|=|{\operatorname{div}_{f}\tilde{h}}|\leq\mathcal{O}(e^{-Af})\quad\text{for all }A<1.$
Since $(M,g)$ has volume growth of order $r^{\frac{n+1}{2}}$, this implies
(3.81) $\int_{M}|Z|^{2}e^{f}<\infty.$
Hence, taking $R\to\infty$ in (3.79) proves the claim that $Z=0$.
It follows that for any compactly supported smooth vector field $Y$,
(3.82) $\int_{M}\langle\tilde{h},\operatorname{div}_{f}^{*}Y\rangle e^{f}=\int_{M}\langle\operatorname{div}_{f}\tilde{h},Y\rangle e^{f}=0.$
We use this with $Y=-2\eta_{R}\tilde{W}$, where $\eta_{R}$ is the same cutoff as above. Recall that $\tilde{h}=\mathcal{L}_{\tilde{W}}g=-2\operatorname{div}_{f}^{*}\tilde{W}$. Then by (3.82),
(3.83) 00 0 $=-2\int_{M}\langle\tilde{h},\operatorname{div}_{f}^{*}(\eta_{R}\tilde{W})\rangle e^{f}=-2\int_{M}\langle\tilde{h},\eta_{R}\operatorname{div}_{f}^{*}\tilde{W}+d\eta_{R}\ast\tilde{W}\rangle e^{f}$
(3.84) $\geq\int_{M}\eta_{R}|\tilde{h}|^{2}e^{f}-C\int_{M}|\tilde{h}||\nabla\eta_{R}||\tilde{W}|e^{f}.$
Since $\nabla\eta_{R}$ is supported on $\{R\leq f\leq 2R\}$ with $|\nabla\eta_{R}|\leq\frac{C}{R}$, and $|\tilde{W}|\leq\mathcal{O}(r^{\frac{1}{2}})$, it follows that
(3.85) $\int_{M}\eta_{R}|\tilde{h}|^{2}e^{f}$ $\leq C\int_{\{R<f<2R\}}|\tilde{h}|e^{f}$
(3.86) $\leq C\mathrm{Vol}(\{R<f<2R\})^{\frac{3}{4}}\left(\int_{M}|\tilde{h}|^{4}e^{4f}\right)^{\frac{1}{4}}$
where the last line uses Hölder’s inequality. By (3.75) and the fact that $(M,g)$ has volume growth of order $r^{\frac{n+1}{2}}$, this gives
(3.87) $\int_{M}\eta_{R}|\tilde{h}|^{2}e^{f}\leq CR^{\frac{3(n+1)}{8}}e^{-\frac{R}{16}}.$
Taking $R\to\infty$, we obtain $\tilde{h}=0$ and this completes the proof. ∎
4. Global symmetry and cohomogeneity one structure
Next, we establish some global symmetry principles. For a vector field $X$ on a complete Riemannian manifold $(M,g)$, let $\Phi^{X}$ be its time 1 flow. Since $(M,g)$ is complete, $\Phi^{X}$ is a well-defined isometry of $(M,g)$ for any Killing field $X$. Let $\mathcal{I}(M,g)$ be the isometry group of $(M,g)$, and $\mathcal{I}^{0}(M,g)$ its identity component. These are finite-dimensional Lie groups by the Myers–Steenrod theorem.
Lemma 4.1.
Let $(M^{n},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS. Let $\{W_{i}\}_{i=1}^{d}$, $d=\dim\mathcal{I}(\mathcal{S},\bar{g}_{\infty})$, be the set of Killing fields on $(M,g)$ obtained from Corollary 3.2. If
(4.1) $d\geq\dim\mathcal{I}(\mathcal{S},g_{\rho})$
for some sufficiently large $\rho>0$, then the collection of isometries
(4.2) $\tilde{G}=\{\Phi^{X}\in\mathcal{I}(M,g):X\in\mathrm{span}_{\mathbb{R}}\{W_{i}\}_{i=1}^{d}\}$
is a Lie subgroup of $\mathcal{I}^{0}(M,g)$. Moreover, restriction to the level set $\{f=\rho\}$ defines a Lie group isomorphism $\tilde{G}\to\mathcal{I}^{0}(\mathcal{S},g_{\rho})$. We therefore have a smooth action of $\mathcal{I}^{0}(\mathcal{S},g_{\rho})$ on $(M,g)$ by isometries.
Proof.
Let $\{W_{i}\}_{i=1}^{d}$ be the Killing fields on $(M,g)$ from Corollary 3.2. Then $\{W_{i}\}_{i=1}^{d}$ restricts to a set of $d$ linearly independent Killing fields on $(\mathcal{S},g_{\rho})$. By (4.1), the restrictions form a basis for the Lie algebra of $\mathcal{I}(\mathcal{S},g_{\rho})$. Therefore, the time 1 flows of vector fields in their span generate $\mathcal{I}^{0}(\mathcal{S},g_{\rho})$:
(4.3) $\tilde{G}_{\rho}:=\{\Phi^{X}|_{\{f=\rho\}}:X\in\mathrm{span}_{\mathbb{R}}\{W_{i}\}_{i=1}^{d}\}=\mathcal{I}^{0}(\mathcal{S},g_{\rho}).$
Let $\tilde{G}$ be defined as in (4.2). We check that $\tilde{G}$ is closed under composition; the other requirements for being a subgroup of $\mathcal{I}(M,g)$ are trivial to check. Let $X,Y\in\mathrm{span}_{\mathbb{R}}\{W_{i}\}_{i=1}^{d}$, and consider the composition $\Phi^{X}\circ\Phi^{Y}$. Since $X$ and $Y$ are tangential to the level sets of $f$, it follows from (4.3) that
(4.4) $(\Phi^{X}\circ\Phi^{Y})|_{\{f=\rho\}}=\Phi^{X}|_{\{f=\rho\}}\circ\Phi^{Y}|_{\{f=\rho\}}=\Phi^{Z}|_{\{f=\rho\}}$
for some $Z\in\mathrm{span}_{\mathbb{R}}\{W_{i}\}_{i=1}^{d}$. Then for any $x\in\{f=\rho\}$, we have
(4.5) $(\Phi^{X}\circ\Phi^{Y})(x)$ $=\Phi^{Z}(x),$
(4.6) $D_{x}(\Phi^{X}\circ\Phi^{Y})|_{T_{x}\{f=\rho\}}$ $=D_{x}\Phi^{Z}|_{T_{x}\{f=\rho\}}.$
Moreover, since $[W_{i},\nabla f]=0$ for any $i$, we have $[X,\nabla f]=0$. Thus, the flows of $X$ and $\nabla f$ commute, so $D\Phi^{X}(\nabla f)=\nabla f$. The same holds for $Y$ and $Z$, so
(4.7) $D_{x}(\Phi^{X}\circ\Phi^{Y})(\nabla f)$ $=D_{x}\Phi^{Z}(\nabla f).$
Since $\Phi^{X}\circ\Phi^{Y}$ and $\Phi^{Z}$ are isometries of $(M,g)$, and they agree to first order at the point $x$ by (4.5), (4.6) and (4.7), it follows that
(4.8) $\Phi^{X}\circ\Phi^{Y}=\Phi^{Z}\quad\text{on }(M,g).$
Hence, $\tilde{G}$ is a subgroup of $\mathcal{I}(M,g)$.
The map $\Phi^{(\cdot)}$, by virtue of being the Lie exponential map, is a diffeomorphism from a neighborhood of the origin in the space of Killing fields on $(M,g)$ to a neighborhood of the identity in $\mathcal{I}(M,g)$. Since $\tilde{G}$ is the image of the restriction of $\Phi^{(\cdot)}$ to the subspace of Killing fields on $(M,g)$ spanned by the $W_{i}$’s, it follows that a neighborhood of the identity in $\tilde{G}$ is a smooth submanifold of $\mathcal{I}(M,g)$. So $\tilde{G}$ is a Lie subgroup of $\mathcal{I}(M,g)$.
Finally, the restriction homomorphism $\tilde{G}\to\tilde{G}_{\rho}$ is clearly smooth and surjective. Injectivity follows from the same argument as above. That is, we check that if two transformations in $\tilde{G}$ agree on $\{f=\rho\}$, then they must agree to first order at $x\in\{f=\rho\}$, and hence coincide everywhere. Thus, $\tilde{G}\cong\tilde{G}_{\rho}=\mathcal{I}^{0}(\mathcal{S},g_{\rho})$. ∎
In applications, the number of Killing fields on $(\mathcal{S},g_{\rho})$ is known. This allows us to narrow down the possibilities for $g_{\rho}$ so that (4.1) can be checked explicitly. If $g_{\rho}$ is a homogeneous metric, then the action of $G$ on $(M,g)$ will be of cohomogeneity one, i.e. $M/G$ is one-dimensional:
Proposition 4.2.
Let $(M^{n},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS. Suppose that for some sufficiently large $\rho>0$, we have
  1. (i)
    $\dim\mathcal{I}(\mathcal{S},\bar{g}_{\infty})\geq\dim\mathcal{I}(\mathcal{S},g_{\rho})$, and
  2. (ii)
    There is a compact, connected Lie group $G$ acting smoothly, transitively, and isometrically on $(\mathcal{S},g_{\rho})$ with isotropy subgroup $H$.
Then $G$ acts smoothly on $(M,g)$ by isometries, and the following hold:
  1. (a)
    The orbit space $M/G$ is homeomorphic to $[0,1)$. We denote the projection map by $\pi:M\to M/G$.
  2. (b)
    $f$ is constant on each orbit.
  3. (c)
    The principal orbits, i.e. $\pi^{-1}(s)$ for any $s>0$, are diffeomorphic to $\mathcal{S}\cong G/H$. The isotropy subgroup of each point in a principal orbit is $H$.
  4. (d)
    The singular orbit $X=\pi^{-1}(0)$ is diffeomorphic to $G/K$, where $K$ is a closed subgroup of $G$ such that $H\subset K$ and $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$.
  5. (e)
    $M$ is diffeomorphic to the total space of the vector bundle $E=G\times_{K}\mathbb{R}^{\ell+1}$ over $G/K\cong X$, where $K$ acts via $k\cdot(g,v)=(gk^{-1},kv)$, and $kv$ is the radial extension of the action of $K$ on $K/H\cong\mathbb{S}^{\ell}\hookrightarrow\mathbb{R}^{\ell+1}$.
  6. (f)
    Every sphere bundle of $E$ is diffeomorphic to $\mathcal{S}$. Thus $M\setminus X$ is diffeomorphic to $(0,\infty)\times\mathcal{S}$, and under this identification the metric $g$ becomes
    (4.9) $g=ds^{2}+g(s),$
    where $\{g(s)\}_{s>0}$ is a smooth family of $G$-invariant metrics on $G/H\cong\mathcal{S}$.
Proof.
Let $\{W_{i}\}_{i=1}^{d}$ be the Killing fields from Corollary 3.2, where $d=\dim\mathcal{I}(\mathcal{S},\bar{g}_{\infty})$. By condition (i) and Lemma 4.1, we have
(4.10) $\mathcal{I}^{0}(\mathcal{S},g_{\rho})\cong\{\Phi^{X}\in\mathcal{I}(M,g):X\in\mathrm{span}_{\mathbb{R}}\{W_{i}\}_{i=1}^{d}\}$
with the isomorphism from right to left given by restriction to the level set $\{f=\rho\}$. By condition (ii), there is a Lie group homomorphism $G\to\mathcal{I}^{0}(\mathcal{S},g_{\rho})$ whose image acts transitively on $(\mathcal{S},g_{\rho})$. In view of the identification (4.10), it follows that $G$ acts isometrically on $(M,g)$, and $f$ is constant on each orbit because $W_{i}$ is orthogonal to $\nabla f$. Moreover, restricting the action to $(\mathcal{S},g_{\rho})$ recovers the transitive action of $G$ from condition (ii). Hence, $(\mathcal{S},g_{\rho})$ is a codimension 1 orbit of the total action with isotropy subgroup $H$. By Mostert [mostert], the existence of a codimension 1 orbit implies that the orbit space $M/G$ is homeomorphic to one of
(4.11) $\mathbb{S}^{1},\quad[0,1],\quad[0,1),\quad(0,1).$
As $\pi:M\to M/G$ is continuous and $M$ is noncompact, we can rule out $\mathbb{S}^{1}$ and $[0,1]$. We can also rule out $(0,1)$, as by [mostert] this would imply $M$ is diffeomorphic to $(0,1)\times\mathcal{S}$, contradicting Theorem 2.4. Hence, $M/G$ is homeomorphic to $[0,1)$.
Thus, we have proved (a) and (b). The other assertions are standard facts about manifolds admitting an isometric action of cohomogeneity one with orbit space homeomorphic to $[0,1)$; see for instance [alexandrino-bettiol]*§6.3 or [hoelscher, mostert]. ∎
5. Uniqueness of the Bryant soliton
5.1. Proof of Theorem 1.4
We now prove the uniqueness of the Bryant soliton, exactly as in [brendle2013rotational, brendle2014rotational, zz].
Proof of Theorem 1.4.
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$. Then $(\mathbb{S}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. Thus, Corollary 3.2 gives $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{n(n-1)/2}$ on $\{f=\rho\}$ for all large $\rho$. It follows that these level sets have constant sectional curvature. The conclusion follows from Proposition 5.2 below. ∎
5.2. Roundness near infinity implies global roundness
In [brendle2013rotational, brendle2014rotational, zz], once rotational symmetry of all sufficiently far level sets of $f$ is established, global rotational symmetry is said to follow easily. We provide the details for completeness, and also because we need a slight generalization of this fact (Proposition 5.2) later.
Lemma 5.1.
Let $G/H$ be one of the following homogeneous spaces:
  • $G/H=SO(n)/SO(n-1)$ (diffeomorphic to $\mathbb{S}^{n-1}$);
  • $G/H=SO(n)/S(O(n-1)\times O(1))$ (diffeomorphic to $\mathbb{RP}^{n-1}$);
  • $G/H=SU(2)/\Gamma$, where $\Gamma\subset SU(2)$ is a finite subgroup.
Then $G/H$ has a unique $G$-invariant metric with constant sectional curvature 1.
Proof.
When $G/H=\mathbb{S}^{n-1}$, this is standard and follows from Schur’s lemma (see e.g. [kerin-wraith]). The same argument works for $G/H=\mathbb{RP}^{n-1}$. For the remaining case, we recall a standard fact about $SU(2)$-invariant metrics: if $g$ is an $SU(2)$-invariant metric on $SU(2)/\Gamma$, with universal Riemannian covering $(SU(2),\tilde{g})$, then there exist positive numbers $\lambda,\mu,\nu>0$ such that
(5.1) $\tilde{g}=\lambda^{2}\sigma_{1}\otimes\sigma_{1}+\mu^{2}\sigma_{2}\otimes\sigma_{2}+\nu^{2}\sigma_{3}\otimes\sigma_{3},$
where $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is the dual coframe of a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
(5.2) $[X_{1},X_{2}]=2X_{3},\quad[X_{2},X_{3}]=2X_{1},\quad[X_{3},X_{1}]=2X_{2}.$
Moreover, $g$ is the quotient of $\tilde{g}$ by the right action of $\Gamma$. Now if $g$ has constant sectional curvature 1, then so does $\tilde{g}$, which forces $\lambda=\mu=\nu=1$. Hence, $\tilde{g}$ is uniquely determined, and so is $g$. ∎
Proposition 5.2.
Let $(M^{n},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}\cong G/H$ is one of
  • $\mathcal{S}=\mathbb{S}^{n-1}$, with $G=SO(n)$ and $H=SO(n-1)$;
  • $\mathcal{S}=\mathbb{RP}^{n-1}$, with $G=SO(n)$ and $H=S(O(n-1)\times O(1))$;
  • $\mathcal{S}=SU(2)/\Gamma$, with $G=SU(2)$ and $H=\Gamma$ a finite subgroup.
If all far enough level sets of $f$ have constant sectional curvature, then $\mathcal{S}=\mathbb{S}^{n-1}$ and $(M,g)$ is the Bryant soliton.
Proof.
A classical theorem of de Rham [derham] says that diffeomorphic spherical space forms are isometric. Thus, all far enough level sets of $f$ are isometric to multiples of $(\mathcal{S},\bar{g}_{\infty})$. This allows us to apply Proposition 4.2. Hence, there is a submanifold $X\subset M$ not homeomorphic to $\mathcal{S}$, such that $(M\setminus X,g)$ can be identified as
(5.3) $(M\setminus X,g)=\left((0,\infty)\times\mathcal{S},ds^{2}+g(s)\right),$
where $\{g(s)\}_{s>0}$ is a family of $G$-invariant metrics on a fixed copy of $G/H\cong\mathcal{S}$. From the description of $(M,g)$ as a vector bundle over $X$, the metric spaces $(\mathcal{S},g(s))$ converge in the Hausdorff sense as $s\to 0$ to $X$ with the metric stucture from $g$. Moreover, $f$ is a function of $s$, and by Lemma 2.1, $f\sim s$ at infinity.
Since all far enough level sets of $f$ have constant sectional curvature, it follows from Lemma 5.1 that
(5.4) $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}\quad\text{for all large }s,$
say for $s\in(\rho_{0},\infty)$, where $w$ is a smooth positive function and $\bar{g}_{\mathcal{S}}$ is the unique $G$-invariant metric with constant sectional curvature 1 on $\mathcal{S}$. Importantly, the $g(s)$ are literally multiples of the fixed metric $\bar{g}_{\mathcal{S}}$; there is no twisting by $s$-dependent diffeomorphisms of $\mathcal{S}$. Thus, the steady GRS equation $\operatorname{Ric}=\mathrm{Hess}_{f}$ applied to (5.3) reduces to the same ODE system as the Bryant soliton, which is
(5.5) $f^{\prime\prime}$ $=-(n-1)\frac{w^{\prime\prime}}{w},$
(5.6) $ww^{\prime\prime}$ $=-w^{\prime}f^{\prime}+(n-2)(1-(w^{\prime})^{2}).$
(See for instance [chow2023ricci]*§6, noting the difference in sign convention for $f$.) Our steady GRS $(M,g,f)$ therefore gives a solution $(w,f)$ to this system on $(\rho_{0},\infty)$. Let $(a,\infty)$, $-\infty\leq a<\rho_{0}$, be the maximal interval on which this solution extends while maintaining $w>0$. The metric $\hat{g}=ds^{2}+w(s)^{2}\bar{g}_{\mathcal{S}}$ on $(a,\infty)\times\mathcal{S}$, being a gradient Ricci soliton, is real-analytic [ivey-analytic]. Likewise, $g$ is real-analytic on $(0,\infty)\times\mathcal{S}$. We have $g=\hat{g}$ on $(\rho_{0},\infty)\times\mathcal{S}$. Hence, unique continuation results (e.g. [kobayashi-nomizu-1]*Lemma §VI.6.3’) apply to give
(5.7) $g=\hat{g}=ds^{2}+w(s)^{2}\bar{g}_{\mathcal{S}}\quad\text{on }(\max\{a,0\},\infty)\times\mathcal{S}.$
If $a<0$, then $w(s)$ has a positive limit as $s\to 0$, so $(\mathcal{S},g(s))$ converges in the Hausdorff sense as $s\to 0$ to a limiting space homeomorphic to $\mathcal{S}$. This is a contradiction since $X$ is not homeomorphic to $\mathcal{S}$. Thus $a\geq 0$. We further claim that
(5.8) $\lim_{s\to a^{+}}w(s)=0.$
To see this, rewrite (5.5) and (5.6) as a system of the form $f^{\prime\prime}=F(f^{\prime},w^{\prime},w)$ and $w^{\prime\prime}=G(f^{\prime},w^{\prime},w)$, where $F$ and $G$ are functions defined wherever $w>0$. Using these expressions and local existence theorems for ODEs, we deduce that at least one of the following is met as $s\to a^{+}$:
  1. (i)
    $f^{\prime}$ becomes unbounded;
  2. (ii)
    $w^{\prime}$ becomes unbounded;
  3. (iii)
    $w$ becomes unbounded;
  4. (iv)
    $w\to 0$.
Now, (i) cannot happen because $|\nabla f|=|f^{\prime}|$ is bounded, and neither can (iii) for obvious geometric reasons. Also, $|f^{\prime\prime}|\leq|{\operatorname{Ric}}|$ is uniformly bounded, so $|w^{\prime\prime}|=\left|\frac{wf^{\prime\prime}}{n-1}\right|$ is bounded near $a$; hence $w^{\prime}$ is also bounded near $a$ and (ii) cannot happen. So $w\to 0$ as claimed.
By (5.7) and (5.8), the metrics $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}$ degenerate to a point as $s\to a^{+}$. If $a>0$, this gives a contradiction in view of (5.3). Hence $a=0$, and $X$ is a point. In order for $(M,g)$ to be smooth, we require that $\mathcal{S}=\mathbb{S}^{n-1}$. In this case, since $(M,g,f)$ is a complete steady GRS by assumption, and is globally rotationally symmetric, it must be the Bryant soliton. ∎
5.3. Proof of Theorem 1.6
This is almost identical to the proof of Theorem 1.4.
Proof of Theorem 1.6.
Let $(M^{n},g,f)$ be an asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS of dimension $n\geq 3$. Then $(\mathbb{RP}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. Exactly as in the proof of Theorem 1.4 (see §5.1), we deduce that each far enough level set of $f$ has constant sectional curvature. The desired contradiction follows from Proposition 5.2. ∎
6. Uniqueness of Appleton’s solitons
In this section we prove Theorem 1.8.
6.1. Two basic facts
We first state two general facts that will be repeatedly used. For groups $H\subseteq G$, let $C_{G}(H)$ and $N_{G}(H)$ be the centralizer and normalizer of $H$ in $G$ respectively. As with $\mathcal{I}$ and $\mathcal{I}^{0}$, we denote identity components by a superscript 00, e.g. $C_{G}^{0}(H)$, $N_{G}^{0}(H)$.
Lemma 6.1.
Let $G$ be a Lie group and $H$ a discrete subgroup. Then $N_{G}^{0}(H)=C_{G}^{0}(H)$.
Proof.
Since $C_{G}(H)\subseteq N_{G}(H)$, we have $C_{G}^{0}(H)\subseteq N_{G}^{0}(H)$. Conversely, let $g\in N_{G}^{0}(H)$. There is a smooth curve $\gamma:[0,1]\to N_{G}^{0}(H)$ such that $\gamma(0)$ is the identity and $\gamma(1)=g$. Since $\gamma(t)\in N_{G}^{0}(H)$, we have
(6.1) $\gamma(t)h\gamma(t)^{-1}\in H\quad\text{for all }t\in[0,1]\text{ and }h\in H.$
For each fixed $h\in H$, the mapping $t\mapsto\gamma(t)h\gamma(t)^{-1}$ is smooth, and sends 00 to $h$. Since $H$ is discrete, it follows that $\gamma(t)h\gamma(t)^{-1}=h$ for all $t\in[0,1]$. Hence, $\gamma(t)\in C_{G}(H)$ for each $t\in[0,1]$. In particular, $g=\gamma(1)\in C_{G}(H)$. This implies that $N_{G}^{0}(H)\subseteq C_{G}(H)$, which further implies $N_{G}^{0}(H)\subseteq C_{G}^{0}(H)$. ∎
The next lemma is standard:
Lemma 6.2.
Let $(\tilde{M},\tilde{g})$ be a Riemannian manifold and $\Gamma\subset\mathcal{I}(\tilde{M},\tilde{g})$ a finite subgroup of isometries acting freely. Then the quotient manifold $(M,g)$ has isometry group
(6.2) $\mathcal{I}(M,g)$ $=N_{\mathcal{I}(\tilde{M},\tilde{g})}(\Gamma)/\Gamma.$
6.2. Isometries of Berger metrics
Recall some notations from §1.2:
  • For each $m,k\geq 1$, $L_{m,k}$ is the smooth manifold obtained by quotienting $\mathbb{S}^{2m+1}$ by the $\mathbb{Z}_{k}$ action which rotates the Hopf fibers.
  • For $a,b>0$, the metric $g_{a,b}$ on $L_{m,k}$ is obtained by quotienting the Berger metric
    (6.3) $\tilde{g}_{a,b}=a^{2}\sigma\otimes\sigma+b^{2}\pi^{*}g_{\mathbb{CP}^{m}}$
    on $\mathbb{S}^{2m+1}$ by the $\mathbb{Z}_{k}$ action. The metric $g_{a,b}$ has constant sectional curvature $\kappa^{-2}$ if and only if $a=b=\kappa$.
The next two lemmas lay out well-known facts about the isometry groups of $\tilde{g}_{a,b}$ and $g_{a,b}$, as well as their homogeneous presentations. Recall that a metric is homogeneous if its isometry group acts transitively, or equivalently if it admits a set of Killing fields spanning the tangent space at each point.
Lemma 6.3.
Let $m,k\geq 1$ and $a,b>0$.
  1. (a)
    $U(m+1)$ embeds into $\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$ as a transitive subgroup of isometries, with isotropy subgroup $U(m)$ (embedded in $U(m+1)$ as $A\mapsto\mathrm{diag}(A,1)$).
  2. (b)
    $(L_{m,k},g_{a,b})$ is the quotient of $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$ by a central subgroup $\mathbb{Z}_{k}\subset U(1)\cong Z(U(m+1))$.
  3. (c)
    $U(m+1)$ acts transitively and isometrically on $(L_{m,k},g_{a,b})$ with isotropy subgroup $U(m)\times\mathbb{Z}_{k}$ (block-diagonally embedded in $U(m+1)$).
Proof.
Parts (a) and (b) are standard, and (c) follows because the subgroup of $U(m+1)$ generated by the upper-left $U(m)$ and the central $\mathbb{Z}_{k}$ is the indicated copy of $U(m)\times\mathbb{Z}_{k}$. ∎
Lemma 6.4.
Let $m\geq 1$ and $a,b>0$. Then
(6.4) $\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$$\cong\begin{cases}O(2m+2)&\text{if }a=b,\\ U(m+1)\rtimes\mathbb{Z}_{2}&\text{if }a\neq b.\end{cases}$
If additionally $k\geq 3$, then
(6.5) $\mathcal{I}(L_{m,k},g_{a,b})$ $\cong(U(m+1)\rtimes\mathbb{Z}_{2})/\mathbb{Z}_{k}.$
Proof.
If $a=b$, then $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$ is a round sphere with isometry group $O(2m+2)$. If $a\neq b$, it is well-known (e.g. [torralbo-urbano]*§2) that $\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})=\hat{\mathcal{I}}$ where
(6.6) ^=$\hat{\mathcal{I}}=\{A\in O(2m+2):AJ=\pm JA\},$
and $J\in O(2m+2)$ is the matrix associated with multiplication by $i$ in $\mathbb{C}^{m+1}\cong\mathbb{R}^{2m+2}$. One checks that $\hat{\mathcal{I}}$ is the disjoint union
(6.7) ^$\hat{\mathcal{I}}=U(m+1)\sqcup KU(m+1),$
where $U(m+1)$ is considered as embedded in $O(2m+2)$, and $K\in O(2m+2)$ is the matrix which sends $(z_{0},\ldots,z_{m})\in\mathbb{C}^{m+1}\cong\mathbb{R}^{2m+2}$ to $(\bar{z}_{0},\ldots,\bar{z}_{m})$. Then
(6.8) ^$\hat{\mathcal{I}}\cong U(m+1)\rtimes\mathbb{Z}_{2}.$
By Lemma 6.2 and Lemma 6.3(b), we have $\mathcal{I}(L_{m,k},g_{a,b})\cong N_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}(\mathbb{Z}_{k})/\mathbb{Z}_{k}$, where $\mathbb{Z}_{k}$ sits centrally in $U(m+1)\subset\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$. Thus, to prove (6.5) we must show that $N_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}(\mathbb{Z}_{k})\cong U(m+1)\rtimes\mathbb{Z}_{2}$. We divide into two cases again.
Suppose $a\neq b$. With $K\in O(2m+2)$ as above, we see that $KAK^{-1}=\bar{A}$ for all $A\in U(m+1)\subset O(2m+2)$. Thus, $K$ normalizes the indicated central copy of $\mathbb{Z}_{k}$. Using (6.7) and (6.8), it follows that
(6.9) $N_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}(\mathbb{Z}_{k})$ $=U(m+1)\sqcup KU(m+1)\cong U(m+1)\rtimes\mathbb{Z}_{2},$
as required.
On the other hand, suppose $a=b$, so that $\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})=O(2m+2)$. The matrix $Z=\exp(\frac{2\pi}{k}J)$ is a generator of the indicated copy of $\mathbb{Z}_{k}$. Thus,
(6.10) $Z^{\pm 1}=\cos\left(\frac{2\pi}{k}\right)I\pm\sin\left(\frac{2\pi}{k}\right)J,$
where $I$ is the identity matrix of size $22m+$ 2. If $A\in N_{O(2m+2)}(\mathbb{Z}_{k})$, then $AZA^{-1}=Z^{\alpha}$ for some $\alpha\in\mathbb{Z}$, i.e.
(6.11) $A\exp\left(\frac{2\pi}{k}J\right)A^{-1}=\exp\left(\frac{2\pi\alpha}{k}J\right).$
The left matrix has eigenvalues $e^{\pm\frac{2\pi i}{k}}$, each with multiplicity $m+1$. The right matrix has eigenvalues $e^{\pm\frac{2\pi i\alpha}{k}}$, each with multiplicity $m+1$. For these eigenvalues to match, we need $\alpha\equiv\pm 1$ mod $k$. Then $AZA^{-1}=Z^{\pm 1}$. Writing these two equations out in terms of (6.10), then simplifying, we get
(6.12) $\sin\left(\frac{2\pi}{k}\right)(AJ-JA)=0\quad\text{or}\quad\sin\left(\frac{2\pi}{k}\right)(AJ+JA)=0.$
Since $k\geq 3$, we have $\sin(\frac{2\pi}{k})\neq 0$, so $AJ=\pm JA$ and $A\in\hat{\mathcal{I}}$. Thus, $N_{O(2m+2)}(\mathbb{Z}_{k})\subseteq\hat{\mathcal{I}}\cong U(m+1)\rtimes\mathbb{Z}_{2}$. Reversing these arguments gives ^$\hat{\mathcal{I}}\subseteq N_{O(2m+2)}(\mathbb{Z}_{k})$. ∎
Next, we characterize the metrics $\tilde{g}_{a,b}$ and $g_{a,b}$ in terms of the dimensions of their isometry groups. For $\tilde{g}_{a,b}$, this will follow from Ziller’s [ziller-homog] classification of triples $(G,\ell,\tilde{g})$ where $G$ is a compact, connected Lie group acting effectively, transitively, and isometrically on $(\mathbb{S}^{\ell},\tilde{g})$.
Lemma 6.5.
Let $\tilde{g}$ be a homogeneous metric on $\mathbb{S}^{2m+1}$ such that $\dim\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g})\geq(m+1)^{2}$. Then $\tilde{g}$ is isometric to $\tilde{g}_{a,b}$ for some $a,b>0$.
Proof.
By assumption, $G=\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g})$ acts effectively, transitively, and isometrically on $(\mathbb{S}^{2m+1},\tilde{g})$, and $\dim G\geq(m+1)^{2}$. By Ziller’s classification [ziller-homog], we check that the only possibilities for $G$ and $\tilde{g}$ with $\dim G\geq(m+1)^{2}$ are
  • $G\cong SO(2m+2)$ and $\tilde{g}$ is a round metric.
  • $G\cong U(m+1)$ and $\tilde{g}$ is a Berger metric $\tilde{g}_{a,b}$ for some $a,b>0$.
  • If $m=3$, $G\cong\mathrm{Spin}(7)$ and $\tilde{g}$ is a round metric.
In all cases, $(\mathbb{S}^{2m+1},\tilde{g})$ is isometric to $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$ for some $a,b>0$. ∎
Lemma 6.6.
Let $m,k\geq 1$ and $a,b>0$ with $a\neq b$. Suppose that $\mathbb{Z}_{k}$ acts freely and isometrically on $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$, and that the quotient manifold has an isometry group of dimension at least $(m+1)^{2}$. Then the quotient manifold is isometric to $(L_{m,k},g_{a,b})$.
Proof.
By assumption, there is an embedding
(6.13) $\iota:\mathbb{Z}_{k}\hookrightarrow\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$
so that the quotient has an isometry group of dimension at least $(m+1)^{2}$; thus by Lemma 6.2,
(6.14) $\dim N_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}^{0}(\iota(\mathbb{Z}_{k}))\geq(m+1)^{2}.$
But Lemma 6.4 gives $\dim\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})=(m+1)^{2}$, so $N_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}^{0}(\iota(\mathbb{Z}_{k}))=\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$. Lemma 6.1 implies $C_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}^{0}(\iota(\mathbb{Z}_{k}))=\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$, and hence
(6.15) $\iota(\mathbb{Z}_{k})$ $\subset C_{\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})}(\mathcal{I}^{0}(\mathbb{S}^{2m+1},\tilde{g}_{a,b}))$
(6.16) $\cong C_{U(m+1)\rtimes\mathbb{Z}_{2}}(U(m+1))$
(6.17) $=Z(U(m+1))\cong U(1),$
where the first equality in the last line can be checked directly. Hence, $\iota(\mathbb{Z}_{k})$ sits in the central $U(1)$ of $U(m+1)$, so by part (b) of Lemma 6.3, the quotient manifold is isometric to $(L_{m,k},g_{a,b})$. ∎
Corollary 6.7.
Let $m,k\geq 1$, and suppose $g$ is a homogeneous metric on $L_{m,k}$ such that $\dim\mathcal{I}(L_{m,k},g)\geq(m+1)^{2}$. Then $g$ is isometric to $g_{a,b}$ for some $a,b>0$.
Proof.
Since diffeomorphic spherical space forms are isometric [derham], if $g$ has constant sectional curvature $\kappa^{-2}>0$, then it is isometric to $g_{\kappa,\kappa}$.
We now suppose that $g$ has non-constant sectional curvature. The metric $g$ on $L_{m,k}$ lifts to a non-constantly curved homogeneous metric $\tilde{g}$ on the universal cover $\mathbb{S}^{2m+1}$. As each isometry also lifts, we have $\dim\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g})\geq(m+1)^{2}$. By Lemma 6.5, $(\mathbb{S}^{2m+1},\tilde{g})$ is isometric to $(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$ for some $a,b>0$. We have $a\neq b$ since $\tilde{g}$ has non-constant sectional curvature. By construction, $(L_{m,k},g)$ is the quotient of $(\mathbb{S}^{2m+1},\tilde{g})$ by a free, isometric action of $\mathbb{Z}_{k}$. Lemma 6.6 shows that $(L_{m,k},g)$ is isometric to $(L_{m,k},g_{a,b})$. ∎
6.3. Proof of Theorem 1.8
The basic idea is similar to the proofs of Theorems 1.4 and 1.6, but rather than showing that the level sets of $f$ have constant sectional curvature, we can only show they are isometric to $(L_{m,k},g_{a,b})$ using Corollary 6.7. We will combine this with the next few results to obtain the desired conclusions.
Lemma 6.8.
Let $m\geq 1$, and consider $U(m)$ as being embedded in $U(m+1)$ via $A\mapsto\mathrm{diag}(A,1)$. Let $K$ be a closed subgroup of $U(m+1)$ containing $U(m)$. Then one of the following holds:
  • $K=U(m)\times C$ for some subgroup $C$ of $U(1)$;
  • $K=U(m+1)$;
  • If $m=1$, there is the additional possibility that $K=(U(1)\times U(1))\rtimes\mathbb{Z}_{2}$.
Proof.
The identity component $K^{0}$ of $K$ acts canonically on $\mathbb{C}^{m+1}$, and splits $\mathbb{C}^{m+1}$ into irreducible modules which are orthogonal with respect to the standard Hermitian metric. As $K^{0}$ contains the upper-left $U(m)$, there are two possibilities:
  1. (i)
    The decomposition is $\mathbb{C}^{m}\oplus\mathbb{C}^{1}$.
  2. (ii)
    The decomposition is $\mathbb{C}^{m+1}$, i.e. $K^{0}$ acts irreducibly.
Case (i): in this case, $K^{0}$ contains the upper-left $U(m)$ and is a subgroup of
(6.18) $\{g\in U(m+1):g(\mathbb{C}^{m})=\mathbb{C}^{m},\,g(\mathbb{C}^{1})=\mathbb{C}^{1}\}=U(m)\times U(1).$
Thus $K^{0}=U(m)\times C$ for some subgroup $C$ of $U(1)$. As $K^{0}$ is connected, the only possibilities are $K^{0}=U(m)$ or $K^{0}=U(m)\times U(1)$. Since $K^{0}$ is a normal subgroup of $K$, we have $K\subseteq N_{U(m+1)}(K^{0})$. Using this, we find the possibilities for $K$:
  • If $K^{0}=U(m)$, then $N_{U(m+1)}(K^{0})=U(m)\times U(1)$ so $K$ is of the form $U(m)\times C$.
  • If $K^{0}=U(m)\times U(1)$ and $m\geq 2$, then $N_{U(m+1)}(K^{0})=K^{0}$, so $K=K^{0}$.
  • If $K^{0}=U(m)\times U(1)$ and $m=1$, then $N_{U(m+1)}(K^{0})=K^{0}\rtimes\mathbb{Z}_{2}$, so either $K=K^{0}$ or $K=K^{0}\rtimes\mathbb{Z}_{2}$.
Case (ii): since $K^{0}$ contains $U(m)$, its Lie algebra $\mathfrak{k}\subseteq\mathfrak{u}(m+1)$ contains $\mathfrak{u}(m)$, embedded into the upper-left block. Since $K^{0}$ is connected and acts irreducibly on $\mathbb{C}^{m+1}$, the induced representation of $\mathfrak{k}$ on $\mathbb{C}^{m+1}$ is also irreducible. Hence, there exists $X\in\mathfrak{k}$ with nonzero off-diagonal entries with respect to the vector space decomposition $\mathbb{C}^{m+1}=\mathbb{C}^{m}\oplus\mathbb{C}^{1}$, i.e.
(6.19) $X=\begin{bmatrix}A&u\\ -u^{*}&\alpha\end{bmatrix}$
where $A\in\mathfrak{u}(m)$, $u\in\mathbb{C}^{m\times 1}$, and $\alpha\in i\mathbb{R}$, with $u\neq 0$. Using this and the fact that $\mathfrak{k}$ contains $\mathfrak{u}(m)$, it can be checked by computing commutators that $\mathfrak{k}=\mathfrak{u}(m+1)$. We omit the details. It follows that $K^{0}=U(m+1)$, and $K=U(m+1)$. ∎
Corollary 6.9.
Let $m\geq 1$ and $k\geq 2$, and let $U(m)$ be embedded in $U(m+1)$ via $A\mapsto\mathrm{diag}(A,1)$. Let $K$ be a closed subgroup of $U(m+1)$ containing $U(m)\times\mathbb{Z}_{k}$, and assume $K/(U(m)\times\mathbb{Z}_{k})$ is diffeomorphic to $\mathbb{S}^{\ell}$ for some $\ell\geq 0$. Then $K=U(m)\times\mathbb{Z}_{2k}$ or $K=U(m)\times U(1)$.
By Lemma 6.3(c), there is a diffeomorphism $U(m+1)/(U(m)\times\mathbb{Z}_{k})\cong L_{m,k}$. We fix such an identification for the next lemma, which is an analog of Lemma 5.1.
Lemma 6.10.
Any $U(m+1)$-invariant metric on $U(m+1)/(U(m)\times\mathbb{Z}_{k})\cong L_{m,k}$ is equal to $g_{a,b}$ for some $a,b>0$.
Proof.
Such metrics bijectively correspond to $\mathrm{Ad}(U(m))$-invariant inner products on $\mathfrak{p}$, where $\mathfrak{u}(m+1)=\mathfrak{u}(m)\oplus\mathfrak{p}$ is a fixed $\mathrm{Ad}(U(m))$-invariant decomposition. It is easy to see that $\mathfrak{p}$ further decomposes into irreducible $\mathrm{Ad}(U(m))$-modules $\mathfrak{p}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$, with $\dim_{\mathbb{R}}\mathfrak{p}_{1}=1$ and $\dim_{\mathbb{R}}\mathfrak{p}_{2}=2m$. By Schur’s lemma, all $\mathrm{Ad}(U(m))$-invariant inner products on $\mathfrak{p}$ are of the form
$a^{2}\langle\cdot,\cdot\rangle_{\mathfrak{p}_{1}}+b^{2}\langle\cdot,\cdot\rangle_{\mathfrak{p}_{2}},$
where $a,b>0$ and $\langle\cdot,\cdot\rangle_{\mathfrak{p}_{i}}$ is the Euclidean inner product on $\mathfrak{p}_{i}$. Tracing back identifications, this corresponds to the metric $g_{a,b}$ on $L_{m,k}$. The lemma follows. ∎
Finally, we can conclude by proving the uniqueness of Appleton’s solitons.
Proof of Theorem 1.8.
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$. Then $(L_{m,k},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $(m+1)^{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{(m+1)^{2}}$ by Lemma 6.4. Since $(L_{m,k},\bar{g}_{\infty})$ is homogeneous, the Killing fields span the tangent space of $L_{m,k}$ at each point.
Thus, Corollary 3.2 gives $(m+1)^{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{(m+1)^{2}}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $L_{m,k}$ and are equipped with a homogeneous metric with isometry group of dimension at least $(m+1)^{2}$. By Corollary 6.7, each of these level sets is isometric to some $(L_{m,k},g_{a,b})$. Therefore:
  • By Lemma 6.4, $\dim\mathcal{I}(L_{m,k},g_{a,b})=(m+1)^{2}=\dim\mathcal{I}(L_{m,k},\bar{g}_{\infty})$.
  • By Lemma 6.3(c), each far enough level set admits a smooth, transitive action of $U(m+1)$ by isometries, and the isotropy subgroup is $U(m)\times\mathbb{Z}_{k}$.
Thus, we may apply Proposition 4.2 with $G=U(m+1)$ and $H=U(m)\times\mathbb{Z}_{k}$ to get the following.
  1. (i)
    Outside a submanifold $X$, we can identify
    (6.20) $(M\setminus X,g)=\left((0,\infty)\times G/H,ds^{2}+g(s)\right),$
    where $\{g(s)\}_{s>0}$ are $G$-invariant metrics on a fixed copy of $G/H\cong L_{m,k}$.
  2. (ii)
    $f$ is a function of $s$, and $f$ is constant on $X$.
  3. (iii)
    There is a closed subgroup $K$ of $G$ containing $H$ such that $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$, and $X\cong G/K$.
  4. (iv)
    From the description of $(M,g)$ as a vector bundle over $X$, the metric spaces $(G/H,g(s))$ converge in the Hausdorff sense to $X$ as $s\to 0$.
From (i) and Lemma 6.10, we have $g(s)=g_{a(s),b(s)}$ for some positive functions $a,b:(0,\infty)\to\mathbb{R}$. Combining this with (ii) and the computations in [appleton]*Appendix A, the triple $(f,a,b)$ satisfies (1.7)–(1.9). Moreover, by asymptotic $L_{m,k}$-cylindricity,
(6.21) $\lim_{s\to\infty}\frac{a(s)}{b(s)}=1.$
It remains to examine the boundary conditions. By (iii) and Corollary 6.9, we have either $K=U(m)\times\mathbb{Z}_{2k}$ (giving $X\cong L_{m,2k}$), or $K=U(m)\times U(1)$ (giving $X\cong\mathbb{CP}^{m}$). However, the first option is incompatible with (iv) because it is impossible for $(G/H,g_{a(s),b(s)})$ to converge in the Hausdorff sense to a metric space homeomorphic to $L_{m,2k}$. Hence $K=U(m)\times U(1)$ and $X\cong\mathbb{CP}^{m}$.
For $(G/H,g_{a(s),b(s)})$ to converge in the Hausdorff sense to a metric on $\mathbb{CP}^{m}$, we need $a(s)\to 0$ and $b(s)\to b_{0}>0$ as $s\to 0$. For the metric to close up smoothly at $X$, we adapt the discussion in [verdiani-ziller]*§3 to obtain $a^{\prime}(0)=k$ and $b^{\prime}(0)=0$ as necessary conditions. As $f$ is a function of the radial coordinate $s$ on fibers, we also see by restricting to a single fiber that smoothness requires $f^{\prime}(0)=0$. Finally, by Appleton’s analysis [appleton]*Corollary 3.5 and Lemma 4.2, we need $f^{\prime\prime}(0)>0$ for $g$ to be a complete, non-Ricci flat metric.
With the boundary conditions from the last paragraph, we can now rescale $g$ and translate $f$ to normalize the boundary conditions to (1.10), all the while preserving the system (1.7)–(1.9) and the limiting behavior (6.21). This completes the proof that $(M,g,f)\in\mathcal{A}_{m,k}$. ∎
7. Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs
In this section we prove Theorem 1.10 and Theorem 1.11.
7.1. Homogeneous metrics on $SU(2)/\Gamma$
Throughout this section, we will identify $SU(2)$ as the multiplicative group of unit quaternions
(7.1) $SU(2)=\{a+bi+cj+dk\in\mathbb{H}:a^{2}+b^{2}+c^{2}+d^{2}=1\}.$
Any left-invariant metric $\tilde{g}$ on $SU(2)$ is of the form
(7.2) $\tilde{g}=\lambda^{2}\sigma_{1}\otimes\sigma_{1}+\mu^{2}\sigma_{2}\otimes\sigma_{2}+\nu^{2}\sigma_{3}\otimes\sigma_{3},$
where $\lambda,\mu,\nu>0$ are constants, and $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is the dual coframe of a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
(7.3) $[X_{1},X_{2}]=2X_{3},\quad[X_{2},X_{3}]=2X_{1},\quad[X_{3},X_{1}]=2X_{2}.$
The relations (7.3) are also satisfied by the basis $\{i,j,k\}$ for $\mathfrak{su}(2)=T_{1}SU(2)$. Hence, by pulling back by an automorphism of $SU(2)$, we may assume $\{X_{1},X_{2},X_{3}\}$ coincides with $\{i,j,k\}$ at the identity. The isometry groups of left-invariant metrics on $SU(2)$ are well-known (see e.g. [ha-lee]*Theorem 3.8):
Lemma 7.1.
Let $\tilde{g}$ be a left-invariant metric on $SU(2)$ of the form (7.2). The isometry group $\mathcal{I}(SU(2),\tilde{g})$ is given as follows.
  1. (a)
    If $\lambda=\mu=\nu$, then $\mathcal{I}(SU(2),\tilde{g})$ is generated by left multiplications in $SU(2)$, right multiplications in $SU(2)$, and inversion $q\mapsto q^{-1}$ in $SU(2)$. Thus,
    (7.4) $\mathcal{I}(SU(2),\tilde{g})=(SU(2)_{L}\times_{\mathbb{Z}_{2}}SU(2)_{R})\rtimes\mathbb{Z}_{2}.$
  2. (b)
    If exactly two of $\lambda,\mu,\nu$ are equal, then without loss of generality we may assume $\lambda\neq\mu=\nu$. Then $\mathcal{I}(SU(2),\tilde{g})$ is generated by left multiplications in $SU(2)$, right multiplications by $e^{i\theta}$, and the map $q\mapsto j^{-1}qj$. Thus,
    (7.5) $\mathcal{I}(SU(2),\tilde{g})=(SU(2)_{L}\times_{\mathbb{Z}_{2}}U(1)_{R})\rtimes\mathbb{Z}_{2}.$
  3. (c)
    If $\lambda,\mu,\nu$ are distinct, then $\mathcal{I}(SU(2),\tilde{g})$ is generated by left multiplications in $SU(2)$, the map $q\mapsto i^{-1}qi$, and the map $q\mapsto j^{-1}qj$. Thus,
    (7.6) $\mathcal{I}(SU(2),\tilde{g})=SU(2)_{L}\rtimes(\mathbb{Z}_{2})^{2}.$
The finite subgroups of $SU(2)$ up to conjugacy are listed below, alongside their generators and normalizers (see for instance [isoms-3, cmp-su2]). Here, $\mathrm{Pin}(2)=\{e^{i\theta}\}\cup\{je^{i\theta}\}$.
Finite subgroup $\Gamma\subset SU(2)$ $|\Gamma|$ Generators $N_{SU(2)}(\Gamma)$
Cyclic $\mathbb{Z}_{k}$ ($k\geq 1)$ $k$ $e^{2\pi i/k}$ $\mathrm{Pin}(2)$
Binary dihedral $\tilde{D}_{2}$ $8$ $i,j$ $\tilde{O}$
Binary dihedral $\tilde{D}_{k}$ ($k\geq 3$) $4k$ $e^{\pi i/k},j$ $\tilde{D}_{k}$
Binary tetrahedral $\tilde{T}$ $24$ $j$, $e^{\pi i/4}(\frac{1}{\sqrt{2}}i+\frac{1}{\sqrt{2}}j)$ $\tilde{O}$
Binary octahedral $\tilde{O}$ $48$ $\frac{1}{\sqrt{2}}i+\frac{1}{\sqrt{2}}j$, $e^{\pi i/4}(\frac{1}{\sqrt{2}}i+\frac{1}{\sqrt{2}}j)$ $\tilde{O}$
Binary icosahedral $\tilde{I}$ $120$ $j$, $\frac{1}{2}+\frac{1}{2}\sqrt{1-2\cos(\frac{2\pi}{5})}i+\cos(\frac{\pi}{5})j$ $\tilde{I}$
Table 2. Finite subgroups of $SU(2)$ up to conjugacy.
Given a finite subgroup $\Gamma$, we denote by $SU(2)/\Gamma$ the smooth manifold obtained by quotienting $SU(2)$ by the right action of $\Gamma$. For some left-invariant metrics on $SU(2)$, the right action of $\Gamma$ is isometric and the resulting metric on $SU(2)/\Gamma$ is homogeneous. Using Lemma 7.1, we can determine exactly when this happens. We state this for non-cyclic $\Gamma$, which are the only cases relevant for this section.
Corollary 7.2.
Let $\tilde{g}$ be a left-invariant metric on $SU(2)$ of the form (7.2), and let $\Gamma$ be a finite non-cyclic subgroup of $SU(2)$. Suppose one of the following holds:
  1. (i)
    $\Gamma=\tilde{D}_{2}$.
  2. (ii)
    $\Gamma=\tilde{D}_{k}$ for some $k\geq 3$, and at least two of $\lambda,\mu,\nu$ are equal.
  3. (iii)
    $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, and $\lambda=\mu=\nu$.
Then $\tilde{g}$ factors through the right action of $\Gamma$ on $SU(2)$, and the induced metric $g$ on $SU(2)/\Gamma$ is homogeneous with $\dim\mathcal{I}(SU(2)/\Gamma,g)=3$.
Proof.
We sketch the proof for case (i). The other cases are similar.
Let $\Gamma=\tilde{D}_{2}$. By Table 2, $\Gamma$ is generated by $i$ and $j$. In all three cases of Lemma 7.1, right-multiplications by $i$ and $j$ both belong to the isometry group $\mathcal{I}(SU(2),\tilde{g})$. Hence, $\tilde{g}$ factors through the right action of $\Gamma$ on $SU(2)$, which we denote by $\Gamma_{R}$. By Lemma 6.2, the isometry group of the quotient is
(7.7) $\mathcal{I}(SU(2)/\Gamma,g)=N_{\mathcal{I}(SU(2),\tilde{g})}(\Gamma_{R})/\Gamma_{R}.$
In each case of Lemma 7.1, the normalizer $N_{\mathcal{I}(SU(2),\tilde{g})}(\Gamma_{R})$ contains $SU(2)_{L}$ as well as a discrete part coming from the normalizer of $\Gamma$ in $SU(2)$ (see Table 2). Thus, $\mathcal{I}^{0}(SU(2)/\Gamma,g)\cong SU(2)_{L}$ has dimension 3 and acts transitively on $SU(2)/\Gamma$. ∎
We prove a converse of Corollary 7.2:
Lemma 7.3.
Let $\Gamma$ be a finite non-cyclic subgroup of $SU(2)$, and let $g$ be a homogeneous metric on $SU(2)/\Gamma$. Let $(SU(2),\tilde{g})$ be the Riemannian universal cover of $(SU(2)/\Gamma,g)$. Then $\tilde{g}$ is a left-invariant metric of the form (7.2), and one of the following holds:
  • $\Gamma=\tilde{D}_{2}$.
  • $\Gamma=\tilde{D}_{k}$ for some $k\geq 3$, and at least two of $\lambda,\mu,\nu$ are equal.
  • $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, and $\lambda=\mu=\nu$.
Moreover, $(SU(2)/\Gamma,g)$ is isometric to the quotient of $(SU(2),\tilde{g})$ by the right action of $\Gamma$, and $\dim\mathcal{I}(SU(2)/\Gamma,g)=3$.
Proof.
The homogeneous metric $g$ on $SU(2)/\Gamma$ lifts to a homogeneous metric on $\mathbb{S}^{3}$, which by [sekigawa]*Theorem C must be isometric to $SU(2)$ with a left-invariant metric $\tilde{g}$. Hence, $\tilde{g}$ is of the form (7.2), and there is an embedding $\iota:\Gamma\hookrightarrow\mathcal{I}(SU(2),\tilde{g})$ such that $\iota(\Gamma)$ acts freely on $SU(2)$ with quotient space isometric to $(SU(2)/\Gamma,g)$. As $g$ is homogeneous, $\mathcal{I}(SU(2)/\Gamma,g)$ has dimension at least 3, so by Lemma 6.2,
(7.8) $\dim N_{\mathcal{I}(SU(2),\tilde{g})}(\iota(\Gamma))=\dim\mathcal{I}(SU(2)/\Gamma,g)\geq 3.$
We split into cases for the numbers $\lambda,\mu,\nu$ in (7.2).
Case 1: $\lambda,\mu,\nu$ are distinct. Then by (7.8) and Lemma 7.1(iii), $N_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))$ is a connected subgroup of $\mathcal{I}^{0}(SU(2),\tilde{g})\cong SU(2)_{L}$ of dimension $\geq 3$. Combining this with Lemma 6.1 gives $C_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))=N_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))=SU(2)_{L}$. Thus,
(7.9) $\iota(\Gamma)\subset C_{\mathcal{I}(SU(2),\tilde{g})}(SU(2)_{L})\cong\tilde{D}_{2},$
where the last $\tilde{D}_{2}$ consists of right-multiplications by the standard copy of $\tilde{D}_{2}$ (the last isomorphism in (7.9) can be checked directly). As $\Gamma$ is non-cyclic, this is only possible if $\Gamma=\tilde{D}_{2}$, in which case $g$ is isometric to the quotient of $\tilde{g}$ by the right action of $\tilde{D}_{2}$.
Case 2: exactly two of $\lambda,\mu,\nu$ are equal. Without loss of generality, assume $\lambda\neq\mu=\nu$. Then by (7.8) and Lemma 7.1(ii), $N_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))$ is a connected dimension $\geq 3$ subgroup of the group
(7.10) $\mathcal{I}^{0}(SU(2),\tilde{g})=SU(2)_{L}\times_{\mathbb{Z}_{2}}U(1)_{R}\cong U(2).$
Thus, $N_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))\cong H$ where $H\in\{U(2),SU(2)\}$. Lemma 6.1 implies $C_{\mathcal{I}(SU(2),\tilde{g})}^{0}(\iota(\Gamma))\cong H$. Therefore, $\iota(\Gamma)\subset C_{\mathcal{I}(SU(2),\tilde{g})}(H)$. We further divide into two cases.
  • Case 2a: $H=U(2)$. Then similarly to (6.17), one computes
    (7.11) $C_{\mathcal{I}(SU(2),\tilde{g})}(H)\cong C_{U(2)\rtimes\mathbb{Z}_{2}}(U(2))=Z(U(2))\cong U(1).$
    Hence, $\iota(\Gamma)$ embeds as a subgroup of $U(1)$. But this is impossible as $\Gamma$ is non-cyclic by assumption.
  • Case 2b: $H=SU(2)$. Since every copy of $SU(2)$ in $U(2)$ is unique up to conjugation, we may assume $H$ is the preimage of $SU(2)_{L}\times\{1\}$ in $U(2)\cong SU(2)_{L}\times_{\mathbb{Z}_{2}}U(1)_{R}$. Then one checks that
    (7.12) $C_{\mathcal{I}(SU(2),\tilde{g})}(H)=C_{(SU(2)_{L}\times_{\mathbb{Z}_{2}}U(1)_{R})\rtimes\mathbb{Z}_{2}}(SU(2)_{L})$
    is generated by right-multiplications by $e^{i\theta}$ and $j$. Hence, $C_{\mathcal{I}(SU(2),\tilde{g})}(H)\cong\mathrm{Pin}(2)$, whose finite subgroups are either cyclic or binary dihedral. Since $\iota(\Gamma)\subset C_{\mathcal{I}(SU(2),\tilde{g})}(H)$ and $\Gamma$ is not cyclic, it follows that $\Gamma=\tilde{D}_{k}$ for some $k\geq 2$. Then $\iota(\Gamma)$ consists of right-multiplications by the standard copy of $\tilde{D}_{k}$. Hence, $g$ is isometric to the quotient of $\tilde{g}$ by the right action of $\tilde{D}_{k}$.
Case 3: $\lambda=\mu=\nu$. Then $\tilde{g}$ has constant sectional curvature and so does $g$. Since diffeomorphic spherical space forms are isometric [derham], $g$ is isometric to the quotient of $\tilde{g}$ by the right action of $\Gamma$. Here, there is no restriction on $\Gamma$.
In each case, we have identified $(SU(2)/\Gamma,g)$ as the quotient of $(SU(2),\tilde{g})$ by the right action of $\Gamma$. Then $\dim\mathcal{I}(SU(2)/\Gamma,g)=3$ by Corollary 7.2. ∎
7.2. Proof of Theorem 1.10
We now prove the non-existence of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
Proof of Theorem 1.10.
Let $(M^{n},g,f)$ be an asymptotically $SU(2)/\Gamma$-cylindrical steady GRS, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$. By Corollary 7.2, $(SU(2)/\Gamma,\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has 3 linearly independent Killing fields $\{\bar{U}_{1},\bar{U}_{2},\bar{U}_{3}\}$ spanning the tangent space at each point. Thus, Corollary 3.2 gives 3 linearly independent Killing fields $\{W_{1},W_{2},W_{3}\}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $SU(2)/\Gamma$ and are equipped with homogeneous metrics. By Lemma 7.3, the level sets have constant sectional curvature. The desired contradiction follows from Proposition 5.2. ∎
7.3. Uniqueness of $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
Finally, we turn to asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs for $k\geq 2$. There turn out to exist such GRSs, given by $\mathbb{Z}_{2}$-quotients of Appleton solitons in $\mathcal{A}_{1,2k}$. To see this, we use the notation of §1.2. For any $a,b>0$, the manifold $(L_{1,2k},g_{a,b})$ is the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of the subgroup $\mathbb{Z}_{2k}$ generated by $e^{\pi i/k}$. By Lemma 7.1, right multiplication by $j$ belongs to $\mathcal{I}(SU(2),\tilde{g}_{a,b})$. This descends to a fixed-point-free isometric involution of $(L_{1,2k},g_{a,b})$. Hence, the quotient of $(L_{1,2k},g_{a,b})$ by this involution is exactly the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of $\tilde{D}_{k}$. Denote this quotient by $(SU(2)/\tilde{D}_{k},\hat{g}_{a,b})$.
Let $A\in\mathcal{A}_{1,2k}$. Outside an $\mathbb{S}^{2}$ core, $A$ is the Riemannian manifold
(7.13) $\left((0,\infty)\times L_{1,2k},ds^{2}+g_{a(s),b(s)}\right).$
Therefore, the involution from the last paragraph induces a fixed-point-free isometric involution of (7.13), with quotient
(7.14) $\left((0,\infty)\times SU(2)/\tilde{D}_{k},ds^{2}+\hat{g}_{a(s),b(s)}\right).$
In the Hopf fibration $SU(2)\to\mathbb{S}^{2}$, right multiplication by $j$ in $SU(2)$ projects to the antipodal map on the base $\mathbb{S}^{2}$. Hence, the $\mathbb{Z}_{2}$-quotient map $(L_{1,2k},g_{a,b})\to(SU(2)/\tilde{D}_{k},\hat{g}_{a,b})$ degenerates to the quotient map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$ if we send $a\to 0$. It follows that the $\mathbb{Z}_{2}$-quotient map from (7.13) to (7.14) extends globally to $A$, acting as the map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$ on the round $\mathbb{S}^{2}$ core. As the soliton potential $f$ of $A$ depends only on $s$, this $\mathbb{Z}_{2}$-action preserves $f$ and induces a steady GRS denoted $A/\mathbb{Z}_{2}$.
By construction, $A/\mathbb{Z}_{2}$ is diffeomorphic to a 2-plane bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$. This is the total space of the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. Furthermore, since $A$ is asymptotically $L_{1,2k}$-cylindrical, $A/\mathbb{Z}_{2}$ is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical.
Now, we prove the converse: every asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS must be of the form $A/\mathbb{Z}_{2}$.
Proof of Theorem 1.11.
Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$. Following the proof of Theorem 1.10, we deduce that all far enough level sets of $f$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$ and are equipped with homogeneous metrics. Therefore:
  • By Lemma 7.3, the isometry group of each far enough level set has dimension 3, as does the isometry group of $(SU(2)/\tilde{D}_{k},\bar{g}_{\infty})$.
  • Each far enough level set admits a smooth, transitive, isometric action of $SU(2)$ with isotropy subgroup $\tilde{D}_{k}$.
Thus, we may apply Proposition 4.2 with $G=SU(2)$ and $H=\tilde{D}_{k}$ to get the following.
  1. (i)
    Outside a submanifold $X$, we can identify
    (7.15) $(M\setminus X,g)=\left((0,\infty)\times SU(2)/\tilde{D}_{k},ds^{2}+g(s)\right),$
    where $\{g(s)\}_{s>0}$ are homogeneous metrics on a fixed copy of $SU(2)/\tilde{D}_{k}$.
  2. (ii)
    $f$ is a function of $s$, and $f$ is constant on $X$.
  3. (iii)
    There is a closed subgroup $K$ of $SU(2)$ containing $H$ such that $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$, and $X\cong SU(2)/K$.
  4. (iv)
    From the description of $(M,g)$ as a vector bundle over $X$, the metric spaces $(SU(2)/\tilde{D}_{k},g(s))$ converge in the Hausdorff sense to $X$ as $s\to 0$.
By (i) and Lemma 7.3, each $g(s)$ (for $s>0$) is isometric to the right $\tilde{D}_{k}$-quotient of a left-invariant metric $\tilde{g}(s)$ on $SU(2)$. Therefore, as in (7.2), we can write
(7.16) $\tilde{g}(s)=\lambda(s)^{2}\sigma_{1}\otimes\sigma_{1}+\mu(s)^{2}\sigma_{2}\otimes\sigma_{2}+\nu(s)^{2}\sigma_{3}\otimes\sigma_{3},$
where $\lambda,\mu,\nu>0$ are positive functions of $s$.
By (iii) and the classification of closed subgroups of $SU(2)$ (see e.g. [cmp-su2]), we have $K=\tilde{D}_{2k}$ (giving $X\cong SU(2)/\tilde{D}_{2k}$) or $K=\mathrm{Pin}(2)$ (giving $X\cong\mathbb{RP}^{2}$). However, since the covering metrics of $g(s)$ are given by (7.16), the Hausdorff limit of $(SU(2)/\tilde{D}_{k},g(s))$ as $s\to 0$ must be homeomorphic to either $SU(2)/\tilde{D}_{k}$ or a space of lower dimension. This and (iv) rule out $K=\tilde{D}_{2k}$. Thus, $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$. Then $M$ is diffeomorphic to the total space of a vector bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$. This is exactly the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. We have $\pi_{1}(M)=\pi_{1}(\mathbb{RP}^{2})=\mathbb{Z}_{2}$.
A smooth double cover of $M$ is obtained by pulling back the bundle $M$ by the antipodal map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$. This is a 2-plane bundle $\tilde{M}$ over $\mathbb{S}^{2}$ with Euler number ±$\pm 2k$, i.e. the total space of $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$, whose sphere bundles are diffeomorphic to $L_{1,2k}$. Lifting the soliton structure from $M$ to $\tilde{M}\cong\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$ and using (ii), we see that the lifted soliton $(\tilde{M},\tilde{g},\tilde{f})$ is an asymptotically $L_{1,2k}$-cylindrical steady GRS. By Theorem 1.8, we have $(\tilde{M},\tilde{g},\tilde{f})\in\mathcal{A}_{1,2k}$. ∎
Appendix A Asymptotic cylindricity of Appleton’s solitons
In this appendix we prove:
Proposition A.1.
Let $m,k\geq 1$. Then any $(M,g,f)\in\mathcal{A}_{m,k}$ is asymptotically $L_{m,k}$-cylindrical.
Proving this is a matter of assembling various facts from Appleton’s analysis in [appleton]; there it was already shown that any $(M,g,f)\in\mathcal{A}_{m,k}$ has the same asymptotic profile as the Bryant soliton quotiented by $\mathbb{Z}_{k}$. It is more convenient to use the following equivalent definition of asymptotic cylindricity:
Definition A.2.
Let $\mathcal{S}$ be a spherical space form of dimension $n-1$. A steady GRS $(M^{n},g,f)$ is asymptotically $\mathcal{S}$-cylindrical if for any sequence of points $p_{i}\in M$ going to infinity, we have convergence of pointed manifolds
(A.1) $(M,R(p_{i})g,p_{i})\to(\mathcal{S}\times\mathbb{R},\bar{g}_{\mathcal{S}}+dz\otimes dz,p_{\infty})$
in the Cheeger–Gromov sense, without passing to a subsequence. Here, $R$ is the scalar curvature of $g$, $\bar{g}_{\mathcal{S}}$ is a metric on $\mathcal{S}$ with constant sectional curvature normalized to have scalar curvature 1, and $p_{\infty}$ is a point on $\mathcal{S}\times\mathbb{R}$.
Let $m,k\geq 1$, and fix $(M,g,f)\in\mathcal{A}_{m,k}$. Recall from §1.2 that outside the $\mathbb{CP}^{m}$ core, $(M,g)$ is $(0,\infty)\times L_{m,k}$ equipped with a metric of the form
(A.2) $g=ds^{2}+g_{a(s),b(s)},$
where $g_{a(s),b(s)}$ is the $\mathbb{Z}_{k}$-quotient of the Berger metric $\tilde{g}_{a(s),b(s)}$ on $\mathbb{S}^{2m+1}$. Let $Q=a/b$.
Lemma A.3 ([appleton]).
The functions $a$ and $b$ satisfy the ODEs
(A.3) $a^{\prime\prime}$ $=\frac{2mQ^{4}}{a}-2m\frac{(a^{\prime})^{2}}{a}-\left(f^{\prime}-2m\frac{Q^{\prime}}{Q}\right)a^{\prime},$
(A.4) $b^{\prime\prime}$ $=2\frac{m+1-Q^{2}}{b}-2m\frac{(b^{\prime})^{2}}{b}-\left(f^{\prime}+\frac{Q^{\prime}}{Q}\right)b^{\prime}.$
Moreover, after possibly rescaling the soliton $(M,g,f)$, it holds that
(A.5) $\lim_{s\to\infty}Q$ $=1,$
(A.6) $\lim_{s\to\infty}Q^{\prime}$ $=0,$
(A.7) $\lim_{s\to\infty}f^{\prime}$ $=1,$
and
(A.8) $a(s)\sim\sqrt{4ms},\quad b(s)\sim\sqrt{4ms}\quad\text{as }s\to\infty.$
Proof.
The ODEs (A.3) and (A.4) follow from (1.8) and (1.9) respectively, while (A.5) holds by the definition of $\mathcal{A}_{m,k}$. The limit (A.6) is due to [appleton]*Lemma 6.4. The limit $\lim_{s\to\infty}f^{\prime}$ exists and is positive by [appleton]*Lemma 4.2. By scaling $g$, it can be arranged to equal $1$, establishing (A.7). Finally, (A.8) is [appleton]*Corollary 6.7. ∎
The next ODE lemma follows from the proof of [appleton]*Lemma 6.2:
Lemma A.4.
Let $y:[0,\infty)\to\mathbb{R}$ be a solution to the ODE
(A.9) $y^{\prime\prime}=\frac{c_{1}(s)}{2y}-\alpha\frac{(y^{\prime})^{2}}{y}-c_{2}(s)y^{\prime},$
where $\alpha>1$ is a constant, and $c_{i}:[0,\infty)\to\mathbb{R}$, $i=1,2$, are two positive functions converging to positive limits $c_{i}^{*}>0$ as $s\to\infty$. Then the function $z=y^{\alpha+1}$ satisfies
(A.10) $\frac{z^{\prime}}{z^{\frac{\alpha-1}{\alpha+1}}}\sim\frac{\alpha+1}{2}\frac{c_{1}^{*}}{c_{2}^{*}}\quad\text{as }s\to\infty.$
Proof of Proposition A.1.
Let $(M,g,f)\in\mathcal{A}_{m,k}$ with $m,k\geq 1$. Using Lemma A.3, we can apply Lemma A.4 to $a$ and $b$, giving
(A.11) $aa^{\prime}\sim 2m,\quad bb^{\prime}\sim 2m\quad\text{as }s\to\infty.$
Then by (A.8),
(A.12) $a^{\prime}(s)\sim\frac{1}{\sqrt{ms}},\quad b^{\prime}(s)\sim\frac{1}{\sqrt{ms}}\quad\text{as }s\to\infty.$
Inserting this into (A.3) and (A.4), and also using the other limits from Lemma A.3, we find that the leading terms cancel and hence
(A.13) $a^{\prime\prime}(s)=o(s^{-\frac{1}{2}}),\quad b^{\prime\prime}(s)=o(s^{-\frac{1}{2}}).$
Next, we use the formula
(A.14) $R=-2m\frac{a^{\prime\prime}}{a}-4m\frac{b^{\prime\prime}}{b}-4m\frac{a^{\prime}b^{\prime}}{ab}-2m(2m-1)\left(\frac{b^{\prime}}{b}\right)^{2}-2m\frac{a^{2}}{b^{4}}+\frac{2m(2m+2)}{b^{2}}$
(see [appleton]*Appendix A) and the above asymptotics to see that
(A.15) $R\sim\frac{2m+1}{2s}.$
By repeatedly differentiating (A.3) and (A.4), one can obtain for each $k\in\mathbb{N}$
(A.16) $a^{(k)}(s)=o(1),\quad b^{(k)}(s)=o(1).$
For each $i\in\mathbb{N}$, define the functions $a_{i}:\mathbb{R}\to\mathbb{R}$ and $b_{i}:\mathbb{R}\to\mathbb{R}$ by
(A.17) $a_{i}(z)$ $=\sqrt{R(s_{i})}a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),$
(A.18) $b_{i}(z)$ $=\sqrt{R(s_{i})}b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right).$
By (A.15) and (A.16), it holds for each $k\geq 0$ that
(A.19) $a_{i}\to\sqrt{2m(2m+1)},\quad b_{i}\to\sqrt{2m(2m+1)}\quad\text{in }C^{k}_{\mathrm{loc}}(\mathbb{R})$
as $i\to\infty$.
Now, let $p_{i}\in M$ be a sequence of points such that $s_{i}=s(p_{i})\to\infty$. Since the scalar curvature (A.14) is a function of $s$, we have $R(p_{i})=R(s_{i})$. Then for each $i$,
(A.20) $R(p_{i})g$ $=R(s_{i})(ds^{2}+g_{a(s),b(s)})$
(A.21) $=dz^{2}+R(s_{i})g_{a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right)}$
(A.22) $=dz^{2}+g_{a_{i}(z),b_{i}(z)},$
where $z=\sqrt{R(s_{i})}(s-s_{i})$.
The metric on $L_{m,k}$ given by
(A.23) $\bar{g}=g_{\sqrt{2m(2m+1)},\sqrt{2m(2m+1)}}$
has constant sectional curvature, and is normalized to have constant scalar curvature 1. Thus, to obtain the required pointed convergence (A.1), it suffices to check that for all $k\geq 0$ and $\rho>0$, the sequence
(A.24) $\sup_{|z|\leq\rho}\left\lVert g_{a_{i}(z),b_{i}(z)}-\bar{g}\right\rVert_{C^{k}(\bar{g})}$
converges to 00 as $i\to\infty$. This convergence indeed holds because of (A.19). ∎

### v2

@@
 We also prove that any steady gradient Ricci soliton asymptotic to a cylinder over the homogeneous lens space $\mathbb{S}^{2m+1}/\mathbb{Z}_{k}=L_{m,k}$, for $m\geq 1$ and $k\geq 3$, is a noncollapsed Appleton soliton on the complex line bundle $\mathcal{O}(-k)$ over $\mathbb{CP}^{m}$.
-In dimension 4, our results lead to a classification of steady gradient Ricci soliton singularity models on smooth manifolds which possess a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$.
+Specializing to dimension 4, we classify steady gradient Ricci soliton singularity models on smooth orbifolds with tangent flows at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$.
 1. Introduction
-A Ricci soliton is a complete $n$-dimensional Riemannian manifold $(M^{n},g)$ with a complete vector field $X$ such that
-(1.1) $\operatorname{Ric}=\frac{1}{2}\mathcal{L}_{X}g+\lambda g,$
-where $\mathcal{L}$ is the Lie derivative, and $\lambda\in\mathbb{R}$ determines the soliton type: shrinking ($\lambda>0$), steady ($\lambda=0$), or expanding ($\lambda<0$).
-If $X=\nabla f$ for a function $f:M\to\mathbb{R}$, then $(M^{n},g,f)$ is called a gradient Ricci soliton and the equation (1.1) becomes
-(1.2) $\operatorname{Ric}=\mathrm{Hess}_{f}+\lambda g.$
-Throughout this article, we abbreviate GRS = gradient Ricci soliton.
-Ricci solitons are self-similar solutions of the Ricci flow.
-Shrinking and steady Ricci solitons are special instances of ancient solutions, which are solutions of Ricci flow defined for all negative time.
-Ancient solutions arise as blowup limits of Ricci flow at singularities; we call these blowup limits singularity models.
-Classifying Ricci solitons and ancient solutions provides a way to understand singularity models, which is in turn crucial for understanding singularity formation in Ricci flow.
-In dimension 3, Perelman [perelman1] showed that finite-time singularity models of Ricci flow are ancient $\kappa$-solutions: non-flat, $\kappa$-noncollapsed ancient solutions with bounded and nonnegative sectional curvature.
-The classification of ancient $\kappa$-solutions in dimension 3 was completed by Brendle [brendle-acta], building on techniques developed in his work on 3-dimensional steady GRSs [brendle2013rotational] and the foundations laid by many other authors.
-For more information on Ricci solitons and ancient solutions in dimensions $\leq 3$, we refer to the books [chow2023ricci, chowluni] and references therein.
-In dimensions $\geq 4$, ancient $\kappa$-solutions no longer capture all singularity models of Ricci flow.
-In fact, singularity models in general need not enjoy any curvature bound unless one imposes strong positivity conditions on the initial data, such as positive isotropic curvature in dimension 4 [hamilton-pic].
-Yet, all known classification results for Ricci solitons and ancient solutions in dimensions $\geq 4$ impose curvature hypotheses alongside other structural assumptions; see for instance [brendle2014rotational, brendle-naff, zz, deng-zhu, chodosh, chodosh-fong, cds].
-This raises the natural question: can one classify Ricci solitons and/or ancient solutions in dimensions $\geq 4$ without imposing any curvature assumptions?
-In this paper, we answer this question affirmatively for steady GRSs that are asymptotic to cylinders over certain spherical space forms.
-The notion of asymptotic cylindricity is standard and arises naturally in various situations; see e.g.
-[brendle2013rotational, deng-zhu, bamler-3].
+Let $(M^{n},g)$ be a complete $n$-dimensional smooth Riemannian manifold or orbifold, and let $f:M\to\mathbb{R}$ be a smooth function.
+The triple $(M^{n},g,f)$ is called a gradient Ricci soliton (GRS) if
+(1.1) $\operatorname{Ric}=\mathrm{Hess}_{f}+\lambda g,$
+where $\lambda\in\mathbb{R}$ determines the soliton type: shrinking ($\lambda>0$), steady ($\lambda=0$), or expanding ($\lambda<0$).
+GRSs generate self-similar solutions of the Ricci flow.
+Shrinking and steady GRSs are special instances of ancient solutions, which are solutions of Ricci flow defined for all negative time.
+Ancient solutions are important in that they arise as blowup limits of Ricci flow at singularities.
+The classification of GRSs and ancient solutions is therefore of significant interest, and has crucial implications for Ricci flow.
+In dimension 3, Perelman [perelman1] showed that finite-time singularities of Ricci flow are modeled on ancient $\kappa$-solutions: non-flat, $\kappa$-noncollapsed ancient solutions on smooth manifolds with bounded and nonnegative sectional curvature.
+The classification of ancient $\kappa$-solutions in dimension 3 was completed by Brendle [brendle-acta] in the noncompact case and by Brendle–Daskalopoulos–Sesum in the compact case [bds], building on Brendle’s work on 3-dimensional steady GRSs [brendle2013rotational] and the foundations laid by many other authors.
+For more information on GRSs and ancient solutions in dimensions $\leq 3$, we refer to the books [chow2023ricci, chowluni] and references therein.
+In dimensions $\geq 4$, ancient $\kappa$-solutions no longer model all singularities of Ricci flow.
+In fact, blowup limits in general need not enjoy any curvature bound unless one imposes strong positivity conditions on the initial data, such as positive isotropic curvature in dimension 4 [hamilton-pic].
+To complicate matters further, Bamler’s compactness and partial regularity theory [bamler-a, bamler-b, bamler-c] shows that singularity models of Ricci flow in dimensions $\geq 4$ are generally not smooth manifolds.
+Yet, virtually all known classification results for GRSs and ancient solutions in dimensions $\geq 4$ take place on smooth manifolds with curvature hypotheses, alongside other structural assumptions; see for instance [brendle2014rotational, brendle-naff, zz, deng-zhu, chodosh, chodosh-fong, cds].
+In this paper, we classify steady GRSs that are asymptotic to cylinders over certain spherical space forms.
+Our results require no curvature hypotheses.
+In dimension 4, our results also hold for steady GRS singularity models on smooth orbifolds with isolated singularities.
+This is the expected regularity for singularity models of 4-dimensional Ricci flow.
+We use a standard notion of asymptotic cylindricity which arises naturally in various situations (see e.g.
+[brendle2013rotational, deng-zhu, bamler-3]):
 Definition 1.1.
@@
 Let $\mathcal{S}=\mathbb{S}^{n-1}/\Gamma$ be a spherical space form of dimension $n-1$.
-A steady GRS $(M^{n},g,f)$ is said to be asymptotically $\mathcal{S}$-cylindrical if the following holds.
-Let $\Phi_{t}$ be the flow of the vector field $\nabla f$.
-Take an arbitrary sequence $p_{m}$ of marked points going to infinity and consider the rescaled flows
+A steady GRS $(M^{n},g,f)$ on a smooth manifold or orbifold is called asymptotically $\mathcal{S}$-cylindrical if the following holds.
+Let $\Phi_{t}$ be the flow of $\nabla f$.
+Take an arbitrary sequence $p_{m}$ of points going to infinity and consider the rescaled flows
 (1.2) $\hat{g}^{(m)}(t)=r_{m}^{-1}\Phi_{r_{m}t}^{*}g,$
@@
 1.1. Uniqueness of the Bryant soliton
-A famous theorem of Brendle [brendle2013rotational, brendle2014rotational] states that any asymptotically cylindrical steady GRS of dimension $n\geq 3$ with positive sectional curvature must be rotationally symmetric, hence isometric to the Bryant soliton [bryant].
+A famous theorem of Brendle [brendle2013rotational, brendle2014rotational] states that any asymptotically cylindrical steady GRS of dimension $n\geq 3$ on a smooth manifold with positive sectional curvature must be rotationally symmetric, hence isometric to the Bryant soliton [bryant].
 In three dimensions, the hypotheses are met for any non-flat, $\kappa$-noncollapsed steady GRS, and Brendle’s result in this setting resolved a major conjecture left open in Perelman’s first paper [perelman1].
-Ten years later, Zhao and Zhu [zz] generalized Brendle’s theorem in dimensions $\geq 4$ by relaxing the positive sectional curvature condition to the weaker condition $2P<R$, where $P$ is the operator norm of the Riemann curvature acting on symmetric 2-tensors.
-Our first result proves uniqueness of the Bryant soliton in a much broader setting.
-We show that uniqueness holds in any dimension only assuming asymptotic cylindricity, without any curvature hypotheses.
+Ten years later, Zhao and Zhu [zz] generalized Brendle’s theorem in dimensions $\geq 4$ to hold under the weaker curvature condition $2P<R$, where $P$ is the operator norm of the Riemann curvature acting on symmetric 2-tensors.
+Deng [deng-orbifold] has also recently extended Brendle’s theorem to smooth orbifolds with positive sectional curvature.
+We show that the curvature hypothesis in Brendle’s theorem can be entirely dropped.
+Thus, uniqueness of the Bryant soliton holds only assuming asymptotic cylindricity for smooth manifolds in any dimension.
 Theorem 1.4.
-Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$. Then $(M,g,f)$ is the Bryant soliton (i.e. they are isometric up to scale).
-In dimension 4, Deng and Zhu [deng-zhu]*Theorem 1.5 established a sufficient condition for a steady GRS to be asymptotically cylindrical. Combining Theorem 1.4 with their result, we get:
+Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$. Then $(M,g,f)$ is the Bryant soliton (i.e. they are isometric up to scale).
+In dimension 4, Deng and Zhu [deng-zhu]*Theorem 1.5 established conditions for a steady GRS on a smooth manifold to be asymptotically cylindrical.
+They conjectured that such a soliton must be the Bryant soliton.
+Theorem 1.4 confirms this:
 Corollary 1.5.
-Any 4-dimensional $\kappa$-noncollapsed steady GRS with positive Ricci curvature and linear curvature decay must be the Bryant soliton.
-This affirmatively answers a conjecture posed by Deng and Zhu in the same paper.
-Corollary 1.5 is comparable to Brendle’s 4-dimensional result [brendle2014rotational], which reaches the same conclusion assuming positive isotropic curvature instead of ‘positive Ricci curvature and linear curvature decay’.
+Any 4-dimensional $\kappa$-noncollapsed steady GRS on a smooth manifold with positive Ricci curvature and linear curvature decay must be the Bryant soliton.
+Corollary 1.5 is comparable to Brendle’s 4-dimensional result [brendle2014rotational], which reaches the same conclusion under the assumption of positive isotropic curvature.
 Theorem 1.4 is proved using Brendle’s Killing field technique, which upgrades asymptotic symmetries to exact ones then exploits them for rigidity and uniqueness.
@@
 Theorem 1.6.
-There is no asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS for any $n\geq 3$.
+There is no asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS on a smooth manifold for any $n\geq 3$.
 1.2. Uniqueness of Appleton’s noncollapsed solitons
-In [appleton], Appleton discovered a new family of $\kappa$-noncollapsed steady GRSs in each even dimension $\geq 4$, and indicated numerical evidence of these solitons arising as Ricci flow singularity models.
+In [appleton], Appleton discovered a new family of $\kappa$-noncollapsed steady GRSs in each even dimension $\geq 4$, and indicated numerical evidence that these solitons do arise as Ricci flow blowup limits.
 A uniqueness theorem is therefore in order, and this is our next result.
@@
   1. (a)
-    Each $(M,g,f)\in\mathcal{A}_{m,k}$ is $U(m+1)$-invariant, $\kappa$-noncollapsed, and asymptotically $L_{m,k}$-cylindrical (see Appendix A).
+    Each $(M,g,f)\in\mathcal{A}_{m,k}$ is $U(m+1)$-invariant, $\kappa$-noncollapsed, and asymptotically $L_{m,k}$-cylindrical.
   2. (b)
@@
 Theorem 1.8.
-Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$. Then $(M,g,f)\in\mathcal{A}_{m,k}$.
+Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS on a smooth manifold, with $m\geq 1$ and $k\geq 3$. Then $(M,g,f)\in\mathcal{A}_{m,k}$.
 Remark 1.9.
 Appleton also conjectured that each $\mathcal{A}_{m,k}$ contains a unique soliton up to scaling and isometry. We do not address this uniqueness question here.
-1.3. 4D steady GRSs with homogeneous 3-cylindrical tangent flows
+1.3. Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs
 Any 3-dimensional spherical space form that is homogeneous with the round metric (i.e.
-its isometry group acts transitively) is of the form $SU(2)/\Gamma$ for some finite subgroup $\Gamma\subset SU(2)$ acting from the right.
-Every finite subgroup of $SU(2)$ is one of the following up to conjugacy (see e.g.
+its isometry group acts transitively) must be $SU(2)/\Gamma$ for some finite subgroup $\Gamma$ acting from the right.
+The finite subgroups of $SU(2)$ are classified up to conjugacy (see e.g.
 [cmp-su2]):
@@
 We have $L_{1,k}\cong SU(2)/\mathbb{Z}_{k}$.
-Thus, Theorems 1.4, 1.6, and 1.8 together classify the asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical steady GRSs for $k\geq 1$.
-To complete the picture, we classify the asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for all other $\Gamma$.
-For $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, no such solitons exist:
+Thus, Theorems 1.4, 1.6, and 1.8 classify the asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical steady GRSs for $k\geq 1$.
+Using the results of Bamler et al.
+[bamler-3], these theorems generalize to 4-dimensional steady GRSs on smooth orbifolds with isolated singularities which are singularity models in the sense of [bamler-3]*Definition 1.1.
+One simply adds in the appropriate Bryant soliton quotients:
+Corollary 1.10.
+Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities. Assume $(M,g,f)$ is asymptotically $SU(2)/\mathbb{Z}_{k}$-cylindrical.
+  1. (a)
+    If $k=1$, then $(M,g,f)$ is the Bryant soliton.
+  2. (b)
+    If $k=2$, then $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{2}$.
+  3. (c)
+    If $k\geq 3$, then either $(M,g,f)\in\mathcal{A}_{m,k}$, or $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$.
+Meanwhile, for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, we will prove the following uniqueness result:
 Theorem 1.11.
-There is no asymptotically $SU(2)/\Gamma$-cylindrical steady GRS for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
-For $\Gamma=\tilde{D}_{k}$ ($k\geq 2$), the story is different.
-Let $E_{k}$ be the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$.
-The sphere bundles of $E_{k}$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$, and there is a double covering $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$.
-We will see that any Appleton soliton $A\in\mathcal{A}_{1,2k}$ descends through the map $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)\to E_{k}$, giving an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS on $E_{k}$ denoted by $A/\mathbb{Z}_{2}$.
-Moreover, we show that $A/\mathbb{Z}_{2}$ are the only steady GRSs with such asymptotics.
+Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities.
+If $(M,g,f)$ is asymptotically $SU(2)/\Gamma$-cylindrical, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$, then $(M,g,f)$ is the Bryant soliton quotiented by $\Gamma$.
+For the remaining case $\Gamma=\tilde{D}_{k}$ ($k\geq 2$), the story is different.
+Each $A\in\mathcal{A}_{1,2k}$ has a $\mathbb{Z}_{2}$-quotient $A/\mathbb{Z}_{2}$, which is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical and lives on the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$.
+See §7.3 for details.
+We prove that $A/\mathbb{Z}_{2}$ are the only steady GRSs with such asymptotics aside from Bryant soliton quotients.
 Theorem 1.12.
-Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$. Then $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
-Combining the 4-dimensional versions of our results (Theorems 1.4, 1.6, 1.8, 1.10, and 1.11), we arrive at the following classification.
-Corollary 1.12.
-Let $(M^{4},g,f)$ be an asymptotically $\mathcal{S}$-cylindrical steady GRS, where $\mathcal{S}=SU(2)/\Gamma$ is a homogeneous 3-dimensional spherical space form with $\Gamma\subset SU(2)$ finite.
-Then exactly one of the following holds:
-  • $\Gamma=\{1\}$, and $(M,g,f)$ is the Bryant soliton.
-  • $\Gamma=\mathbb{Z}_{k}$ for some $k\geq 3$, and $(M,g,f)\in\mathcal{A}_{1,k}$.
-  • $\Gamma=\tilde{D}_{k}$ for some $k\geq 2$, and $(M,g,f)=A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$.
-We highlight one further consequence of this. Bamler et al. [bamler-3] showed that any tangent flow at infinity of a 4-dimensional steady GRS singularity model is one of
-(1.12) $\mathbb{R}^{4}/\Gamma\text{ (but not }\mathbb{R}^{4}\text{)},\quad(\mathbb{S}^{3}/\Gamma)\times\mathbb{R},\quad\mathbb{S}^{2}\times\mathbb{R}^{2},\quad((\mathbb{S}^{2}\times\mathbb{R})/\mathbb{Z}_{2})\times\mathbb{R},$
-where $\Gamma$ is a finite subgroup of $O(4)$.
-Moreover, if one tangent flow at infinity is $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$, then the soliton is asymptotically $\mathbb{S}^{3}/\Gamma$-cylindrical.
-Hence, Corollary 1.12 classifies all steady GRS singularity models with a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$, where $\Gamma$ is a finite subgroup of $SU(2)$.
+Let $(M^{4},g,f)$ be a steady GRS singularity model on a smooth orbifold with isolated singularities.
+If $(M,g,f)$ is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical for some $k\geq 2$, then $(M,g,f)$ is either $A/\mathbb{Z}_{2}$ for some $A\in\mathcal{A}_{1,2k}$, or it is the Bryant soliton quotiented by $\tilde{D}_{k}$.
 1.4. Summary of results
-For the reader’s convenience, we gather our classification results into Table 1 below.
+Table 1 summarizes our classification results for asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth manifolds.
 $\mathcal{S}$ Asymptotically $\mathcal{S}$-cylindrical steady GRSs Theorem
@@
 $L_{m,k}$ ($m\geq 1$, $k\geq 3$) Appleton solitons in $\mathcal{A}_{m,k}$ 1.8
-$SU(2)/\Gamma$ ($\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$) None 1.10
-$SU(2)/\tilde{D}_{k}$ ($k\geq 2$) $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.11
-Table 1. Classification of asymptotically $\mathcal{S}$-cylindrical steady GRSs for various spherical space forms $\mathcal{S}$.
-We also have Corollary 1.5 and Corollary 1.12 as consequences in dimension 4.
-1.5. Remarks on the orbifold case
-While this paper exclusively works with smooth Riemannian manifolds, a robust singularity theory of Ricci flow in dimensions $\geq 4$ requires objects of lower regularity.
-Indeed, Bamler’s compactness theory [bamler-a, bamler-b, bamler-c] shows that finite-time singularity models of Ricci flow are orbifold Ricci solitons with singular locus of codimension $\geq 4$.
-It is therefore of interest to see if our results generalize to the orbifold setting, and we believe they do.
-Our proofs use local properties like the maximum principle and the analyticity of Ricci solitons, which generalize directly to orbifolds as one can pass to local uniformizing covers and apply results for smooth manifolds there.
-We also use the standard formalism for cohomogeneity one manifolds, of which a version for orbifolds exists [gonzalez-alvaro-phd, borzellino-phd].
-To this end, the only major change occurs in Proposition 4.2(d), where $K/H$ is now allowed to be a spherical space form $\mathbb{S}^{\ell}/\Gamma$.
-Naively applying this change to our proofs, it appears that quotients of the Bryant soliton are the only new possible steady GRSs that arise in the expanded category of smooth Riemannian orbifolds.
-Hence, we conjecture:
-Conjecture 1.13.
-Table 1 correctly classifies asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth Riemannian orbifolds for certain $\mathcal{S}$, provided that one includes the appropriate Bryant soliton quotient as an extra possibility in each row.
-Verifying this conjecture requires checking if global and asymptotic properties of steady GRSs on Riemannian manifolds carry over to orbifolds.
-For instance, we used connectedness at infinity (Theorem 2.4) whose proof does not generalize directly to orbifolds.
-Instead, this property is separately established in [bamler-3]*Appendix A for 4-dimensional steady GRS singularity models with a tangent flow at infinity of the form $(\mathbb{S}^{3}/\Gamma)\times\mathbb{R}$.
-In a recent article [deng-orbifold], Deng proved that any asymptotically $\mathcal{S}$-cylindrical steady GRS on a smooth orbifold with positive sectional curvature must be a quotient of the Bryant soliton.
-Unlike for manifolds, it is impossible to drop the curvature hypothesis (at least entirely), as the solitons in $\mathcal{A}_{m,k}$ would furnish counterexamples.
-The reason our methods cannot apply here is because Deng employs a reduction to manifolds that fails without curvature conditions.
-Specifically, Deng obtains a covering of the orbifold GRS $M$ by an asymptotically cylindrical steady GRS on a manifold $\tilde{M}$ with positive sectional curvature.
-Then Brendle’s uniqueness result [brendle2014rotational] (alternatively Theorem 1.4) applies to $\tilde{M}$.
-However, the existence of such a cover requires the singular locus of $M$ to consist of a single point, which demands at least the positivity of Ricci curvature (see [deng-orbifold]*§4).
+$SU(2)/\Gamma$ ($\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$) None 1.11
+$SU(2)/\tilde{D}_{k}$ ($k\geq 2$) $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.12
+Table 1. Classification of asymptotically $\mathcal{S}$-cylindrical steady GRSs on smooth manifolds for various spherical space forms $\mathcal{S}$.
+$\Gamma\subset SU(2)$ Asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs Theorem
+$\mathbb{Z}_{1}=\{1\}$ Bryant soliton 1.10
+$\mathbb{Z}_{2}$ Bryant$/\mathbb{Z}_{2}$ 1.10
+$\mathbb{Z}_{k}$ ($k\geq 3$) Bryant$/\mathbb{Z}_{k}$, and Appleton solitons in $\mathcal{A}_{1,k}$ 1.10
+$\tilde{T}$, $\tilde{O}$, or $\tilde{I}$ Bryant$/\Gamma$ 1.11
+$\tilde{D}_{k}$ ($k\geq 2$) Bryant$/\tilde{D}_{k}$, and $A/\mathbb{Z}_{2}$ for $A\in\mathcal{A}_{1,2k}$ 1.12
+Table 2. Classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRS singularity models on smooth orbifolds with isolated singularities.
+Table 2 classifies asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs which are singularity models on smooth orbifolds with isolated singularities.
+It also completely classifies such steady GRSs with a tangent flow at infinity of the form $(SU(2)/\Gamma)\times\mathbb{R}$, since this implies asymptotic $SU(2)/\Gamma$-cylindricity by [bamler-3].
 1.5. Outline of the paper
-In §2, we collect relevant facts (due to other authors) about asymptotically $\mathcal{S}$-cylindrical steady GRSs.
+In §2, we collect known facts about asymptotically $\mathcal{S}$-cylindrical steady GRSs.
 In particular, any such soliton $(M,g,f)$ opens up like a paraboloid, and the level sets of $f$ converge smoothly (after a certain rescaling) to a spherical space form $(\mathcal{S},\bar{g}_{\infty})$.
@@
 where $\operatorname{Rm}(h)_{ik}=\operatorname{Rm}_{ijkl}h^{jl}$ is the action of the curvature tensor.
-Brendle and Zhao–Zhu [zz] carried this out under extra curvature hypotheses.
-In contrast, our Liouville theorem (Proposition 3.6) is specific to the tensor $h=\mathcal{L}_{W}g$ but requires no curvature hypotheses.
-The proof of Proposition 3.6 uses parts of Brendle’s original blowdown argument, decay estimates of Dung–Sung [dung-sung] and Zhao–Zhu [zz], tensor identities on Ricci solitons (where $h$ being a Lie derivative is essential), maximum principles, and standard integration-by-parts arguments.
+We establish this for the tensor $h=\mathcal{L}_{W}g$ (Proposition 3.6).
+In contrast to Brendle and Zhao–Zhu [zz], this holds without any extra curvature hypotheses.
 In §4, we use the results of §3 to establish global symmetry principles.
-Under mild conditions, we show that the isometry group of each level set of $f$ acts isometrically on the whole of $(M,g)$, and this action preserves all level sets of $f$ (Lemma 4.1).
-Moreover, if the level sets are homogeneous, then the orbit space is one-dimensional (Proposition 4.2).
-In this case, standard facts about cohomogeneity one manifolds apply to pin down the global structure of $(M,g,f)$.
-In §5, we prove the uniqueness of the Bryant soliton (Theorem 1.4) following [brendle2013rotational, brendle2014rotational, zz].
+We show that the isometry group of each level set of $f$ acts isometrically on the whole of $(M,g)$, and this action preserves $f$ (Lemma 4.1).
+If the level sets of $f$ are homogeneous, then $(M,g)$ admits a cohomogeneity one structure (Proposition 4.2).
+In §5, we prove uniqueness of the Bryant soliton (Theorem 1.4) following [brendle2013rotational, brendle2014rotational, zz].
 Given an asymptotically cylindrical steady GRS $(M^{n},g,f)$, we use the results of §3 to find $\frac{n(n-1)}{2}$ Killing fields on each far enough level set of $f$, which implies these level sets are isometric to round spheres.
 From this, we deduce that $(M,g)$ is globally rotationally symmetric.
-This last step was not detailed by Brendle and Zhao–Zhu; we give its proof for completeness, and also because a slight generalization of it (Proposition 5.2) is needed.
-We also prove the non-existence of asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRSs (Theorem 1.6).
-This proceeds almost identically to the proof of Theorem 1.4, except the last step uses Proposition 5.2 to obtain a contradiction.
+By similar arguments, we prove the non-existence of asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRSs on smooth manifolds (Theorem 1.6), as well as parts (a) and (b) of Corollary 1.10 in the orbifold setting.
 In §6, we prove the uniqueness of Appleton’s solitons (Theorem 1.8).
@@
 The results of §3 induce $(m+1)^{2}$ Killing fields on each far enough level set of $f$.
-Using the classification of homogeneous metrics on spheres [ziller-homog] and some group theory, we deduce that the level sets are isometric to quotiented Berger spheres $(L_{m,k},g_{a,b})$.
+From this, we deduce that the level sets are isometric to quotiented Berger spheres $(L_{m,k},g_{a,b})$.
 The results of §4 imply that $(M,g)$ has a cohomogeneity one isometric action of $U(m+1)$, and is specified outside a submanifold $X$ by Appleton’s ODE system (1.6)–(1.8).
-The cohomogeneity one formalism classifies the possible diffeomorphism types of $X$; in our case, we find that the metric completes smoothly if and only if $X\cong\mathbb{CP}^{m}$ and the boundary conditions (1.10) are met.
-We conclude that $(M,g,f)\in\mathcal{A}_{m,k}$.
-In §7, we finish the classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs (Theorems 1.10 and 1.11).
-Here we also describe the solitons $A/\mathbb{Z}_{2}$ (for $A\in\mathcal{A}_{1,2k}$) in more detail.
-Other than this, the ideas and methods involved are similar to those used in §5 and §6, so we omit an outline for brevity.
-Appendix A contains a proof that Appleton’s solitons in $\mathcal{A}_{m,k}$ are asymptotically $L_{m,k}$-cylindrical.
+If $M$ is a smooth manifold, then we deduce that $X\cong\mathbb{CP}^{m}$ and the boundary conditions (1.9) are met.
+We conclude that $(M,g,f)\in\mathcal{A}_{m,k}$, proving Theorem 1.8.
+In the 4-dimensional orbifold setting, we show that $X$ could also be a point, in which case $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$.
+This establishes part (c) of Corollary 1.10.
+In §7, using similar arguments as in §5 and §6, we finish the classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs on orbifolds (Theorems 1.11 and 1.12).
 Acknowledgements
@@
 2. Geometry of asymptotically $\mathcal{S}$-cylindrical steady GRSs
-Throughout this section, $(M^{n},g,f)$ denotes an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}$ is a spherical space form of dimension $n-1$.
+In this section, we assume $(M^{n},g,f)$ is an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}$ is a spherical space form of dimension $n-1$.
+Throughout this paper, all asymptotically $\mathcal{S}$-cylindrical GRSs are assumed to have the following regularity unless otherwise stated:
+  • If $n\neq 4$, then $M$ is a smooth manifold.
+  • If $n=4$, then $M$ is a smooth manifold, or it is a smooth orbifold with isolated singularities and a singularity model in the sense of [bamler-3]*Definition 1.1.
 2.1. Steady soliton identities
@@
 This is self-adjoint with respect to the weighted volume form $e^{f}\mathrm{dvol}_{g}$. Under the normalization (2.3), one has the soliton identity
-(2.5) $\Delta_{f}f=1$
-which will be repeatedly used throughout.
+(2.5) $\Delta_{f}f=1.$
 2.2. Geometric properties
@@
 Proof.
-By [deng-zhu]*Lemma 6.5, asymptotic $\mathcal{S}$-cylindricity implies that $R=o(1)$ as $r\to\infty$. Then the claim follows from (2.3) and Lemma 2.1. ∎
+By [deng-zhu]*Lemma 6.5, asymptotic $\mathcal{S}$-cylindricity implies that $R=o(1)$ as $r\to\infty$.
+This is also true in the 4-dimensional orbifold case [bamler-3]*Theorem 1.3.
+Then the claim follows from (2.3) and Lemma 2.1.
+∎
 Next, we state the roundness estimates on the level sets of $f$ which were first established by Brendle [brendle2014rotational] and subsequently improved by Zhao–Zhu [zz].
@@
 They carry over verbatim to asymptotically $\mathcal{S}$-cylindrical steady GRSs.
-It is clear that $(M,g)$ does not split off a line. Hence:
-Theorem 2.4 ([mw-smms]*Corollary 1.1).
+Theorem 2.4.
 $M$ has exactly one end.
+Proof.
+For smooth manifolds, this follows from [mw-smms]*Corollary 1.1 since $(M,g)$ clearly does not split off a line. In the 4-dimensional orbifold case, see [bamler-3]*Theorem A.1. ∎
 Corollary 2.5.
@@
 3. Killing fields
-Given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, we will show in this section that every symmetry of the asymptotic cross-section $(\mathcal{S},\bar{g}_{\infty})$ manifests as a genuine symmetry of $(M,g)$.
+Given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, we will show in this section that every continuous symmetry of the asymptotic cross-section $(\mathcal{S},\bar{g}_{\infty})$ manifests as a genuine symmetry of $(M,g)$.
 More precisely, we prove:
@@
 where $U_{i}$ is a vector field on $M$ such that $U_{i}|_{\{f=\rho\}}=\bar{U}_{i}$ for all large $\rho>0$. Using (3.2) and Lemma 2.6, and identifying $\{f=\rho\}\cong\mathcal{S}$, we see that
-(3.3) $\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}$ $\leq\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}-\frac{1}{2(n-2)\rho}g_{\rho}}^{2}+\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\frac{1}{2(n-2)\rho}g_{\rho}}^{2}$
-(3.4) $\leq\mathcal{O}(\rho^{-\frac{1}{2}-\frac{1}{2}\delta})\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}+\mathcal{O}(\rho^{-1}).$
-Hence,
 (3.3) $\sup_{\{f=\rho\}}|W_{i}-\bar{U}_{i}|_{\bar{g}_{\infty}}^{2}=\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}\leq\mathcal{O}(\rho^{-1}).$
@@
 For the rest of this section, we fix an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$.
-The goal is to prove Theorem 3.1.
 First, we pull back Killing fields from $(\mathcal{S},\bar{g}_{\infty})$ to obtain approximate Killing fields on $(M,g)$.
@@
 Since $f$ attains a minimum by Lemma 2.1, we conclude that $\langle\tilde{W},\nabla f\rangle$ vanishes identically.
-From this, and the identification (2.9) of the end of $(M,g)$, one checks (e.g.
-by a computation in coordinates) that $\mathcal{L}_{\tilde{W}}g_{\rho}=\mathcal{L}_{\tilde{W}}g=0$ for each large $\rho$.
+From this, and the identification (2.9) of the end of $(M,g)$, one checks that $\mathcal{L}_{\tilde{W}}g_{\rho}=\mathcal{L}_{\tilde{W}}g=0$ for each large $\rho$.
 Thus $\tilde{W}$ is a Killing field on $(\mathcal{S},g_{\rho})$.
@@
 Next, we establish some global symmetry principles.
-For a vector field $X$ on a complete Riemannian manifold $(M,g)$, let $\Phi^{X}$ be its time 1 flow.
-Since $(M,g)$ is complete, $\Phi^{X}$ is a well-defined isometry of $(M,g)$ for any Killing field $X$.
+For a vector field $X$ on $M$, let $\Phi^{X}$ be its time 1 flow.
+If $(M,g)$ is complete, then $\Phi^{X}$ is a well-defined isometry of $(M,g)$ for any Killing field $X$.
 Let $\mathcal{I}(M,g)$ be the isometry group of $(M,g)$, and $\mathcal{I}^{0}(M,g)$ its identity component.
-These are finite-dimensional Lie groups by the Myers–Steenrod theorem.
+These are finite-dimensional Lie groups by the Myers–Steenrod theorem (see [isom-orb]*Theorem 1 for the orbifold version).
 Lemma 4.1.
@@
   4. (d)
-    The singular orbit $X=\pi^{-1}(0)$ is diffeomorphic to $G/K$, where $K$ is a closed subgroup of $G$ such that $H\subset K$ and $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$.
+    The singular orbit $X=\pi^{-1}(0)$ is diffeomorphic to $G/K$, where $K$ is a closed subgroup of $G$ such that $H\subset K$ and $K/H\cong\mathcal{S}^{\prime}$ for some spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$.
+Moreover, $M$ is a smooth manifold if and only if $\mathcal{S}^{\prime}$ is a sphere.
   5. (e)
-    $M$ is diffeomorphic to the total space of the vector bundle $E=G\times_{K}\mathbb{R}^{\ell+1}$ over $G/K\cong X$, where $K$ acts via $k\cdot(g,v)=(gk^{-1},kv)$, and $kv$ is the radial extension of the action of $K$ on $K/H\cong\mathbb{S}^{\ell}\hookrightarrow\mathbb{R}^{\ell+1}$.
+    $M$ is diffeomorphic to the orbifiber bundle $E=G\times_{K}\mathrm{Cone}(\mathcal{S}^{\prime})$ over $G/K\cong X$, where $K$ acts via $k\cdot(g,v)=(gk^{-1},kv)$, and $kv$ is the radial extension of the action of $K$ on $K/H\cong\mathcal{S^{\prime}}$.
   6. (f)
@@
 (4.11) $\mathbb{S}^{1},\quad[0,1],\quad[0,1),\quad(0,1).$
-As $\pi:M\to M/G$ is continuous and $M$ is noncompact, we can rule out $\mathbb{S}^{1}$ and $[0,1]$.
+(This is also true on Riemannian orbifolds, as is implied by the results in [isom-alexandrov]*§2 which apply in this setting.) As $M$ is noncompact, we can rule out $\mathbb{S}^{1}$ and $[0,1]$.
 We can also rule out $(0,1)$, as by [mostert] this would imply $M$ is diffeomorphic to $(0,1)\times\mathcal{S}$, contradicting Theorem 2.4.
@@
 Thus, we have proved (a) and (b).
-The other assertions are standard facts about manifolds admitting an isometric action of cohomogeneity one with orbit space homeomorphic to $[0,1)$; see for instance [alexandrino-bettiol]*§6.3 or [hoelscher, mostert].
+The other assertions are well-known facts about manifolds admitting a cohomogeneity one isometric action with orbit space $[0,1)$; see e.g.
+[alexandrino-bettiol]*§6.3 and [hoelscher]*§1.
+For orbifolds, we refer to [gonzalez-alvaro-phd].
 ∎
-5. Uniqueness of the Bryant soliton
-5.1. Proof of Theorem 1.4
-We now prove the uniqueness of the Bryant soliton, exactly as in [brendle2013rotational, brendle2014rotational, zz].
+5. Uniqueness of the Bryant soliton and its $\mathbb{Z}_{2}$-quotient
+5.1. Uniqueness of the Bryant soliton
+We proceed exactly as in [brendle2013rotational, brendle2014rotational, zz].
 Proof of Theorem 1.4.
-Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$.
+Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$.
 Then $(\mathbb{S}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$.
-Thus, Corollary 3.2 gives $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{n(n-1)/2}$ on $\{f=\rho\}$ for all large $\rho$.
+Thus, Corollary 3.2 gives $\frac{n(n-1)}{2}$ linearly independent Killing fields on $\{f=\rho\}$ for all large $\rho$.
 It follows that these level sets have constant sectional curvature.
-The conclusion follows from Proposition 5.2 below.
+The conclusion follows from Proposition 5.2(a) below.
 ∎
+Part (a) of Corollary 1.10 is proved using the same argument, but instead using part (b) of Proposition 5.2 in the last step.
 5.2. Roundness near infinity implies global roundness
@@
 (5.1) $\tilde{g}=\lambda^{2}\sigma_{1}\otimes\sigma_{1}+\mu^{2}\sigma_{2}\otimes\sigma_{2}+\nu^{2}\sigma_{3}\otimes\sigma_{3},$
-where $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is the dual coframe of a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
+where $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is dual to a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
 (5.2) $[X_{1},X_{2}]=2X_{3},\quad[X_{2},X_{3}]=2X_{1},\quad[X_{3},X_{1}]=2X_{2}.$
@@
   • $\mathcal{S}=SU(2)/\Gamma$, with $G=SU(2)$ and $H=\Gamma$ a finite subgroup.
-If all far enough level sets of $f$ have constant sectional curvature, then $\mathcal{S}=\mathbb{S}^{n-1}$ and $(M,g)$ is the Bryant soliton.
+Suppose that all far enough level sets of $f$ have constant sectional curvature. Then we have the following conclusions depending on the regularity of $M$.
+  1. (a)
+    If $M$ is a smooth manifold, then $\mathcal{S}=\mathbb{S}^{n-1}$ and $(M,g,f)$ is the Bryant soliton.
+  2. (b)
+    If $n=4$ and $M$ is a smooth orbifold with isolated singularities, then $(M,g,f)$ is the corresponding quotient of the Bryant soliton.
 Proof.
@@
 where $\{g(s)\}_{s>0}$ is a family of $G$-invariant metrics on a fixed copy of $G/H\cong\mathcal{S}$.
-From the description of $(M,g)$ as a vector bundle over $X$, the metric spaces $(\mathcal{S},g(s))$ converge in the Hausdorff sense as $s\to 0$ to $X$ with the metric stucture from $g$.
+From the description of $(M,g)$ as a bundle over $X$, the metric spaces $(\mathcal{S},g(s))$ converge in the Hausdorff sense as $s\to 0$ to $X$ with the submanifold metric.
 Moreover, $f$ is a function of $s$, and by Lemma 2.1, $f\sim s$ at infinity.
-Since all far enough level sets of $f$ have constant sectional curvature, it follows from Lemma 5.1 that
+Since all far enough level sets of $f$ have constant sectional curvature, Lemma 5.1 gives
 (5.4) $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}\quad\text{for all large }s,$
@@
 (5.6) $ww^{\prime\prime}$ $=-w^{\prime}f^{\prime}+(n-2)(1-(w^{\prime})^{2}).$
-(See for instance [chow2023ricci]*§6, noting the difference in sign convention for $f$.) Our steady GRS $(M,g,f)$ therefore gives a solution $(w,f)$ to this system on $(\rho_{0},\infty)$.
+(See e.g.
+[chow2023ricci]*§6, noting the difference in sign convention for $f$.) Our steady GRS $(M,g,f)$ therefore gives a solution $(w,f)$ to this system on $(\rho_{0},\infty)$.
 Let $(a,\infty)$, $-\infty\leq a<\rho_{0}$, be the maximal interval on which this solution extends while maintaining $w>0$.
@@
 We have $g=\hat{g}$ on $(\rho_{0},\infty)\times\mathcal{S}$.
-Hence, unique continuation results (e.g.
-[kobayashi-nomizu-1]*Lemma §VI.6.3’) apply to give
+Hence, unique continuation (e.g.
+[kobayashi-nomizu-1]*Lemma §VI.6.3’) applies to give
 (5.7) $g=\hat{g}=ds^{2}+w(s)^{2}\bar{g}_{\mathcal{S}}\quad\text{on }(\max\{a,0\},\infty)\times\mathcal{S}.$
-If $a<0$, then $w(s)$ has a positive limit as $s\to 0$, so $(\mathcal{S},g(s))$ converges in the Hausdorff sense as $s\to 0$ to a limiting space homeomorphic to $\mathcal{S}$.
+If $a<0$, then $w(s)$ has a positive limit as $s\to 0$, so $(\mathcal{S},g(s))$ converges in the Hausdorff sense as $s\to 0$ to a space homeomorphic to $\mathcal{S}$.
 This is a contradiction since $X$ is not homeomorphic to $\mathcal{S}$.
@@
 By (5.7) and (5.8), the metrics $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}$ degenerate to a point as $s\to a^{+}$.
-If $a>0$, this gives a contradiction in view of (5.3).
 Hence $a=0$, and $X$ is a point.
-In order for $(M,g)$ to be smooth, we require that $\mathcal{S}=\mathbb{S}^{n-1}$.
-In this case, since $(M,g,f)$ is a complete steady GRS by assumption, and is globally rotationally symmetric, it must be the Bryant soliton.
+By Proposition 4.2(d), $M$ is diffeomorphic to the cone over $\mathcal{S}$.
+To conclude, we divide into two cases:
+  1. (a)
+    If $M$ is a smooth manifold, we must have $\mathcal{S}=\mathbb{S}^{n-1}$.
+In this case, since $(M,g,f)$ is a complete steady GRS by assumption, and is globally rotationally symmetric, it is the Bryant soliton.
+  2. (b)
+    If $M$ is a 4-dimensional smooth orbifold with isolated singularities, then its universal orbifold cover $\tilde{M}$ is a smooth manifold diffeomorphic to $\mathbb{R}^{4}$.
+Lifting the soliton structure gives an asymptotically cylindrical steady GRS $(\tilde{M},\tilde{g},\tilde{f})$.
+By Theorem 1.4, $(\tilde{M},\tilde{g},\tilde{f})$ is the Bryant soliton.
+5.3. Uniqueness of the $\mathbb{Z}_{2}$-quotient of the Bryant soliton
+The next proof is almost identical to the proof of Theorem 1.4.
+Proof of Theorem 1.6.
+Let $(M^{n},g,f)$ be an asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$.
+Then $(\mathbb{RP}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$.
+As in the proof of Theorem 1.4 (see §5.1), we deduce that each far enough level set of $f$ has constant sectional curvature.
+Proposition 5.2 now gives a contradiction.
 ∎
-5.3. Proof of Theorem 1.6
-This is almost identical to the proof of Theorem 1.4.
-Proof of Theorem 1.6.
-Let $(M^{n},g,f)$ be an asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS of dimension $n\geq 3$.
-Then $(\mathbb{RP}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$.
-Exactly as in the proof of Theorem 1.4 (see §5.1), we deduce that each far enough level set of $f$ has constant sectional curvature.
-The desired contradiction follows from Proposition 5.2.
-∎
+The same argument also proves part (b) of Corollary 1.10.
 6. Uniqueness of Appleton’s solitons
-In this section we prove Theorem 1.8.
+In this section we prove Theorem 1.8, as well as part (c) of Corollary 1.10.
 6.1. Two basic facts
-We first state two general facts that will be repeatedly used.
 For groups $H\subseteq G$, let $C_{G}(H)$ and $N_{G}(H)$ be the centralizer and normalizer of $H$ in $G$ respectively.
@@
 $C_{G}^{0}(H)$, $N_{G}^{0}(H)$.
+We first state two general facts that will be repeatedly used.
+Their proofs are rudimentary and therefore omitted.
 Lemma 6.1.
 Let $G$ be a Lie group and $H$ a discrete subgroup. Then $N_{G}^{0}(H)=C_{G}^{0}(H)$.
-Proof.
-Since $C_{G}(H)\subseteq N_{G}(H)$, we have $C_{G}^{0}(H)\subseteq N_{G}^{0}(H)$.
-Conversely, let $g\in N_{G}^{0}(H)$.
-There is a smooth curve $\gamma:[0,1]\to N_{G}^{0}(H)$ such that $\gamma(0)$ is the identity and $\gamma(1)=g$.
-Since $\gamma(t)\in N_{G}^{0}(H)$, we have
-(6.1) $\gamma(t)h\gamma(t)^{-1}\in H\quad\text{for all }t\in[0,1]\text{ and }h\in H.$
-For each fixed $h\in H$, the mapping $t\mapsto\gamma(t)h\gamma(t)^{-1}$ is smooth, and sends 00 to $h$.
-Since $H$ is discrete, it follows that $\gamma(t)h\gamma(t)^{-1}=h$ for all $t\in[0,1]$.
-Hence, $\gamma(t)\in C_{G}(H)$ for each $t\in[0,1]$.
-In particular, $g=\gamma(1)\in C_{G}(H)$.
-This implies that $N_{G}^{0}(H)\subseteq C_{G}(H)$, which further implies $N_{G}^{0}(H)\subseteq C_{G}^{0}(H)$.
-∎
-The next lemma is standard:
+The following is a standard fact about isometry groups of quotient manifolds:
 Lemma 6.2.
-Let $(\tilde{M},\tilde{g})$ be a Riemannian manifold and $\Gamma\subset\mathcal{I}(\tilde{M},\tilde{g})$ a finite subgroup of isometries acting freely.
+Let $(\tilde{M},\tilde{g})$ be a simply-connected Riemannian manifold and $\Gamma\subset\mathcal{I}(\tilde{M},\tilde{g})$ a finite subgroup of isometries acting freely.
 Then the quotient manifold $(M,g)$ has isometry group
@@
 (6.3) $\mathcal{I}(\mathbb{S}^{2m+1},\tilde{g}_{a,b})$$\cong\begin{cases}O(2m+2)&\text{if }a=b,\\ U(m+1)\rtimes\mathbb{Z}_{2}&\text{if }a\neq b.\end{cases}$
-If additionally $k\geq 3$, then
+If additionally $k\geq 3$, then for any $a,b>0$ we have
 (6.4) $\mathcal{I}(L_{m,k},g_{a,b})$ $\cong(U(m+1)\rtimes\mathbb{Z}_{2})/\mathbb{Z}_{k}.$
@@
 ∎
-6.3. Proof of Theorem 1.8
-The basic idea is similar to the proofs of Theorems 1.4 and 1.6, but rather than showing that the level sets of $f$ have constant sectional curvature, we can only show they are isometric to $(L_{m,k},g_{a,b})$ using Corollary 6.7.
+6.3. Uniqueness of Appleton’s solitons
+The proof of Theorem 1.8 is similar to those of Theorems 1.4 and 1.6, but rather than showing that the level sets of $f$ have constant curvature, we can only show they are isometric to $(L_{m,k},g_{a,b})$ using Corollary 6.7.
 We will combine this with the next few results to obtain the desired conclusions.
@@
 Proof of Theorem 1.8.
-Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$.
+Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS on a smooth manifold, where $m\geq 1$ and $k\geq 3$.
 Then $(L_{m,k},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $(m+1)^{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{(m+1)^{2}}$ by Lemma 6.4.
 Since $(L_{m,k},\bar{g}_{\infty})$ is homogeneous, the Killing fields span the tangent space of $L_{m,k}$ at each point.
-Thus, Corollary 3.2 gives $(m+1)^{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{(m+1)^{2}}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point.
+Corollary 3.2 gives $(m+1)^{2}$ linearly independent Killing fields for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point.
 Hence, all far enough level sets of $f$ are diffeomorphic to $L_{m,k}$ and are equipped with a homogeneous metric with isometry group of dimension at least $(m+1)^{2}$.
@@
   • By Lemma 6.3(c), each far enough level set admits a smooth, transitive action of $U(m+1)$ by isometries, and the isotropy subgroup is $U(m)\times\mathbb{Z}_{k}$.
-Thus, we may apply Proposition 4.2 with $G=U(m+1)$ and $H=U(m)\times\mathbb{Z}_{k}$ to get the following.
+Using Proposition 4.2 with $G=U(m+1)$ and $H=U(m)\times\mathbb{Z}_{k}$, we get the following.
   1. (i)
@@
 (6.20) $\lim_{s\to\infty}\frac{a(s)}{b(s)}=1.$
-It remains to examine the boundary conditions.
-By (iii) and Corollary 6.9, we have either $K=U(m)\times\mathbb{Z}_{2k}$ (giving $X\cong L_{m,2k}$), or $K=U(m)\times U(1)$ (giving $X\cong\mathbb{CP}^{m}$).
+It remains to examine the boundary conditions. By (iii) and Corollary 6.9, there are two possibilities for $K$ and $X$:
+  1. (a)
+    $K=U(m)\times\mathbb{Z}_{2k}$ and $X\cong L_{m,2k}$.
+  2. (b)
+    $K=U(m)\times U(1)$ and $X\cong\mathbb{CP}^{m}$.
 However, the first option is incompatible with (iv) because it is impossible for $(G/H,g_{a(s),b(s)})$ to converge in the Hausdorff sense to a metric space homeomorphic to $L_{m,2k}$.
@@
 This completes the proof that $(M,g,f)\in\mathcal{A}_{m,k}$.
+∎
+Proof of part (c) of Corollary 1.10.
+Let $(M^{4},g,f)$ be an asymptotically $L_{1,k}$-cylindrical steady GRS singularity model on a smooth orbifold with isolated singularities, where $k\geq 3$.
+The proof of Theorem 1.8 applies verbatim, though in item (iii) we also allow $K/H$ to be a spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$.
+Using Lemma 6.8, we see that the only new possibility for $(K,X)$ is that $K=U(2)$ and $X$ is a point.
+Then $M$ is diffeomorphic to a cone over $L_{1,k}$.
+We now run the same argument used in the last paragraph of the proof of Proposition 5.2.
+This shows that $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$.
 ∎
@@
 7.2. Proof of Theorem 1.11
-We now prove the non-existence of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
+We now prove the uniqueness of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
 Proof of Theorem 1.11.
-Let $(M^{n},g,f)$ be an asymptotically $SU(2)/\Gamma$-cylindrical steady GRS, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
+Let $(M^{n},g,f)$ be an asymptotically $SU(2)/\Gamma$-cylindrical steady GRS on a smooth orbifold with isolated singularities, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
 By Corollary 7.2, $(SU(2)/\Gamma,\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has 3 linearly independent Killing fields $\{\bar{U}_{1},\bar{U}_{2},\bar{U}_{3}\}$ spanning the tangent space at each point.
-Thus, Corollary 3.2 gives 3 linearly independent Killing fields $\{W_{1},W_{2},W_{3}\}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point.
+Thus, Corollary 3.2 gives 3 linearly independent Killing fields on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point.
 Hence, all far enough level sets of $f$ are diffeomorphic to $SU(2)/\Gamma$ and are equipped with homogeneous metrics.
 By Lemma 7.3, the level sets have constant sectional curvature.
-The desired contradiction follows from Proposition 5.2.
+The result follows from Proposition 5.2.
 ∎
-7.3. Uniqueness of $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
+7.3. $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
 Finally, we turn to asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs for $k\geq 2$.
-There turn out to exist such GRSs, given by $\mathbb{Z}_{2}$-quotients of Appleton solitons in $\mathcal{A}_{1,2k}$.
+There turn out to exist such GRSs on smooth manifolds, given by $\mathbb{Z}_{2}$-quotients of Appleton solitons in $\mathcal{A}_{1,2k}$.
 To see this, we use the notation of §1.2.
-For any $a,b>0$, the manifold $(L_{1,2k},g_{a,b})$ is the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of the subgroup $\mathbb{Z}_{2k}$ generated by $e^{\pi i/k}$.
+For any $a,b>0$, the manifold $(L_{1,2k},g_{a,b})$ is the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action generated by $e^{\pi i/k}$.
 By Lemma 7.1, right multiplication by $j$ belongs to $\mathcal{I}(SU(2),\tilde{g}_{a,b})$.
@@
 Furthermore, since $A$ is asymptotically $L_{1,2k}$-cylindrical, $A/\mathbb{Z}_{2}$ is asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical.
-Now, we prove the converse: every asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS must be of the form $A/\mathbb{Z}_{2}$.
+7.4. Proof of Theorem 1.12
+We prove that $A/\mathbb{Z}_{2}$ are the only asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs in addition to Bryant soliton quotients.
 Proof of Theorem 1.12.
-Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$.
+Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS on a smooth orbifold with isolated singularities, where $k\geq 2$.
 Following the proof of Theorem 1.11, we deduce that all far enough level sets of $f$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$ and are equipped with homogeneous metrics.
@@
   • Each far enough level set admits a smooth, transitive, isometric action of $SU(2)$ with isotropy subgroup $\tilde{D}_{k}$.
-Thus, we may apply Proposition 4.2 with $G=SU(2)$ and $H=\tilde{D}_{k}$ to get the following.
+Then applying Proposition 4.2 with $G=SU(2)$ and $H=\tilde{D}_{k}$ gives the following:
   1. (i)
@@
   3. (iii)
-    There is a closed subgroup $K$ of $SU(2)$ containing $H$ such that $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$, and $X\cong SU(2)/K$.
+    There is a closed subgroup $K$ of $SU(2)$ containing $H$ such that $K/H\cong\mathcal{S}^{\prime}$ for a spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$.
+We have $X\cong SU(2)/K$.
   4. (iv)
@@
 where $\lambda,\mu,\nu>0$ are positive functions of $s$.
-By (iii) and the classification of closed subgroups of $SU(2)$ (see e.g.
-[cmp-su2]), we have $K=\tilde{D}_{2k}$ (giving $X\cong SU(2)/\tilde{D}_{2k}$) or $K=\mathrm{Pin}(2)$ (giving $X\cong\mathbb{RP}^{2}$).
-However, since the covering metrics of $g(s)$ are given by (7.16), the Hausdorff limit of $(SU(2)/\tilde{D}_{k},g(s))$ as $s\to 0$ must be homeomorphic to either $SU(2)/\tilde{D}_{k}$ or a space of lower dimension.
-This and (iv) rule out $K=\tilde{D}_{2k}$.
-Thus, $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$.
-Then $M$ is diffeomorphic to the total space of a vector bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$.
+By (iii) and the classification of closed subgroups of $SU(2)$ (see e.g. [cmp-su2]), the possibilities for $K$ and $X$ are as follows:
+  1. (a)
+    $K=\tilde{D}_{2k}$ and $X\cong SU(2)/\tilde{D}_{2k}$.
+  2. (b)
+    $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$.
+  3. (c)
+    $K=SU(2)$ and $X$ is a point.
+However, since the covering metrics of $g(s)$ are given by (7.16), the Hausdorff limit of $(SU(2)/\tilde{D}_{k},g(s))$ as $s\to 0$ is homeomorphic to either $SU(2)/\tilde{D}_{k}$ or a space of lower dimension.
+This and (iv) rule out $K=\tilde{D}_{2k}$, so we consider only the other two cases:
+Case (b): $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$.
+In this case $M$ is diffeomorphic to the total space of a vector bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$.
 This is exactly the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$.
-We have $\pi_{1}(M)=\pi_{1}(\mathbb{RP}^{2})=\mathbb{Z}_{2}$.
+In particular, $M$ is a manifold with $\pi_{1}(M)=\pi_{1}(\mathbb{RP}^{2})=\mathbb{Z}_{2}$.
 A smooth double cover of $M$ is obtained by pulling back the bundle $M$ by the antipodal map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$.
@@
 By Theorem 1.8, we have $(\tilde{M},\tilde{g},\tilde{f})\in\mathcal{A}_{1,2k}$.
+Case (c): $K=SU(2)$ and $X$ is a point.
+In this case, $M$ is diffeomorphic to a cone over $\mathcal{S}$.
+We now run the same argument used in the last paragraph of the proof of Proposition 5.2.
+This shows that $(M,g,f)$ is the Bryant soliton quotiented by $\tilde{D}_{k}$.
 ∎
-Appendix A Asymptotic cylindricity of Appleton’s solitons
-In this appendix we prove:
-Proposition A.1.
-Let $m,k\geq 1$. Then any $(M,g,f)\in\mathcal{A}_{m,k}$ is asymptotically $L_{m,k}$-cylindrical.
-Proving this is a matter of assembling various facts from Appleton’s analysis in [appleton]; there it was already shown that any $(M,g,f)\in\mathcal{A}_{m,k}$ has the same asymptotic profile as the Bryant soliton quotiented by $\mathbb{Z}_{k}$.
-It is more convenient to use the following equivalent definition of asymptotic cylindricity:
-Definition A.2.
-Let $\mathcal{S}$ be a spherical space form of dimension $n-1$.
-A steady GRS $(M^{n},g,f)$ is asymptotically $\mathcal{S}$-cylindrical if for any sequence of points $p_{i}\in M$ going to infinity, we have convergence of pointed manifolds
-(A.1) $(M,R(p_{i})g,p_{i})\to(\mathcal{S}\times\mathbb{R},\bar{g}_{\mathcal{S}}+dz\otimes dz,p_{\infty})$
-in the Cheeger–Gromov sense, without passing to a subsequence.
-Here, $R$ is the scalar curvature of $g$, $\bar{g}_{\mathcal{S}}$ is a metric on $\mathcal{S}$ with constant sectional curvature normalized to have scalar curvature 1, and $p_{\infty}$ is a point on $\mathcal{S}\times\mathbb{R}$.
-Let $m,k\geq 1$, and fix $(M,g,f)\in\mathcal{A}_{m,k}$. Recall from §1.2 that outside the $\mathbb{CP}^{m}$ core, $(M,g)$ is $(0,\infty)\times L_{m,k}$ equipped with a metric of the form
-(A.2) $g=ds^{2}+g_{a(s),b(s)},$
-where $g_{a(s),b(s)}$ is the $\mathbb{Z}_{k}$-quotient of the Berger metric $\tilde{g}_{a(s),b(s)}$ on $\mathbb{S}^{2m+1}$. Let $Q=a/b$.
-Lemma A.3 ([appleton]).
-The functions $a$ and $b$ satisfy the ODEs
-(A.3) $a^{\prime\prime}$ $=\frac{2mQ^{4}}{a}-2m\frac{(a^{\prime})^{2}}{a}-\left(f^{\prime}-2m\frac{Q^{\prime}}{Q}\right)a^{\prime},$
-(A.4) $b^{\prime\prime}$ $=2\frac{m+1-Q^{2}}{b}-2m\frac{(b^{\prime})^{2}}{b}-\left(f^{\prime}+\frac{Q^{\prime}}{Q}\right)b^{\prime}.$
-Moreover, after possibly rescaling the soliton $(M,g,f)$, it holds that
-(A.5) $\lim_{s\to\infty}Q$ $=1,$
-(A.6) $\lim_{s\to\infty}Q^{\prime}$ $=0,$
-(A.7) $\lim_{s\to\infty}f^{\prime}$ $=1,$
-and
-(A.8) $a(s)\sim\sqrt{4ms},\quad b(s)\sim\sqrt{4ms}\quad\text{as }s\to\infty.$
-Proof.
-The ODEs (A.3) and (A.4) follow from (1.8) and (1.9) respectively, while (A.5) holds by the definition of $\mathcal{A}_{m,k}$.
-The limit (A.6) is due to [appleton]*Lemma 6.4.
-The limit $\lim_{s\to\infty}f^{\prime}$ exists and is positive by [appleton]*Lemma 4.2.
-By scaling $g$, it can be arranged to equal $1$, establishing (A.7).
-Finally, (A.8) is [appleton]*Corollary 6.7.
-∎
-The next ODE lemma follows from the proof of [appleton]*Lemma 6.2:
-Lemma A.4.
-Let $y:[0,\infty)\to\mathbb{R}$ be a solution to the ODE
-(A.9) $y^{\prime\prime}=\frac{c_{1}(s)}{2y}-\alpha\frac{(y^{\prime})^{2}}{y}-c_{2}(s)y^{\prime},$
-where $\alpha>1$ is a constant, and $c_{i}:[0,\infty)\to\mathbb{R}$, $i=1,2$, are two positive functions converging to positive limits $c_{i}^{*}>0$ as $s\to\infty$.
-Then the function $z=y^{\alpha+1}$ satisfies
-(A.10) $\frac{z^{\prime}}{z^{\frac{\alpha-1}{\alpha+1}}}\sim\frac{\alpha+1}{2}\frac{c_{1}^{*}}{c_{2}^{*}}\quad\text{as }s\to\infty.$
-Proof of Proposition A.1.
-Let $(M,g,f)\in\mathcal{A}_{m,k}$ with $m,k\geq 1$. Using Lemma A.3, we can apply Lemma A.4 to $a$ and $b$, giving
-(A.11) $aa^{\prime}\sim 2m,\quad bb^{\prime}\sim 2m\quad\text{as }s\to\infty.$
-Then by (A.8),
-(A.12) $a^{\prime}(s)\sim\frac{1}{\sqrt{ms}},\quad b^{\prime}(s)\sim\frac{1}{\sqrt{ms}}\quad\text{as }s\to\infty.$
-Inserting this into (A.3) and (A.4), and also using the other limits from Lemma A.3, we find that the leading terms cancel and hence
-(A.13) $a^{\prime\prime}(s)=o(s^{-\frac{1}{2}}),\quad b^{\prime\prime}(s)=o(s^{-\frac{1}{2}}).$
-Next, we use the formula
-(A.14) $R=-2m\frac{a^{\prime\prime}}{a}-4m\frac{b^{\prime\prime}}{b}-4m\frac{a^{\prime}b^{\prime}}{ab}-2m(2m-1)\left(\frac{b^{\prime}}{b}\right)^{2}-2m\frac{a^{2}}{b^{4}}+\frac{2m(2m+2)}{b^{2}}$
-(see [appleton]*Appendix A) and the above asymptotics to see that
-(A.15) $R\sim\frac{2m+1}{2s}.$
-By repeatedly differentiating (A.3) and (A.4), one can obtain for each $k\in\mathbb{N}$
-(A.16) $a^{(k)}(s)=o(1),\quad b^{(k)}(s)=o(1).$
-For each $i\in\mathbb{N}$, define the functions $a_{i}:\mathbb{R}\to\mathbb{R}$ and $b_{i}:\mathbb{R}\to\mathbb{R}$ by
-(A.17) $a_{i}(z)$ $=\sqrt{R(s_{i})}a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),$
-(A.18) $b_{i}(z)$ $=\sqrt{R(s_{i})}b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right).$
-By (A.15) and (A.16), it holds for each $k\geq 0$ that
-(A.19) $a_{i}\to\sqrt{2m(2m+1)},\quad b_{i}\to\sqrt{2m(2m+1)}\quad\text{in }C^{k}_{\mathrm{loc}}(\mathbb{R})$
-as $i\to\infty$.
-Now, let $p_{i}\in M$ be a sequence of points such that $s_{i}=s(p_{i})\to\infty$. Since the scalar curvature (A.14) is a function of $s$, we have $R(p_{i})=R(s_{i})$. Then for each $i$,
-(A.20) $R(p_{i})g$ $=R(s_{i})(ds^{2}+g_{a(s),b(s)})$
-(A.21) $=dz^{2}+R(s_{i})g_{a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right)}$
-(A.22) $=dz^{2}+g_{a_{i}(z),b_{i}(z)},$
-where $z=\sqrt{R(s_{i})}(s-s_{i})$.
-The metric on $L_{m,k}$ given by
-(A.23) $\bar{g}=g_{\sqrt{2m(2m+1)},\sqrt{2m(2m+1)}}$
-has constant sectional curvature, and is normalized to have constant scalar curvature 1.
-Thus, to obtain the required pointed convergence (A.1), it suffices to check that for all $k\geq 0$ and $\rho>0$, the sequence
-(A.24) $\sup_{|z|\leq\rho}\left\lVert g_{a_{i}(z),b_{i}(z)}-\bar{g}\right\rVert_{C^{k}(\bar{g})}$
-converges to 00 as $i\to\infty$. This convergence indeed holds because of (A.19). ∎